### v1

# A Statistical Framework for Understanding Causal Effects that Vary by Treatment Initiation Time in EHR-based Studies

Abstract
Comparative effectiveness studies using electronic health records (EHR) consider data from patients who could “enter” the study cohort at any point during an interval that spans many years in calendar time. Unlike treatments in tightly controlled trials, real-world treatments can evolve over calendar time, especially if comparators include standard of care, or procedures where techniques may improve. Efforts to assess whether treatment efficacy itself is changing are complicated by changing patient populations, with potential covariate shift in key effect modifiers. In this work, we propose a statistical framework to estimate calendar-time specific average treatment effects and describe both how and why effects vary across treatment initiation time in EHR-based studies. Our approach projects doubly robust, time-specific treatment effect estimates onto candidate marginal structural models and uses a model selection procedure to best describe how effects vary by treatment initiation time. We further introduce a novel summary metric, based on standardization analysis, to quantify the role of covariate shift in explaining observed effect changes and disentangle changes in treatment effects from changes in the patient population receiving treatment. Extensive simulations using EHR data from Kaiser Permanente are used to validate the utility of the framework, which we apply to study changes in relative weight loss following two bariatric surgical interventions versus no surgery among patients with severe obesity between 2005–2011.
1 Introduction
Electronic health record (EHR) databases provide a useful data source for comparative effectiveness studies, allowing for larger sample sizes from patient populations which are more diverse than those typical of clinical trials (Casey2016EHR). In EHR-based studies, study entry periods frequently extend over a decade or more (arterburn2020, fisher2018association, coleman2016longterm, Harrington2024, haneuse2025, o2013long), perhaps because (a) treatment initiation is uncommon, (b) long-term outcomes are of interest, and/or (c) extending data collection further back in time requires little additional effort. For example, in a large EHR-based study, arterburn2020 examined long-term weight loss outcomes of patients undergoing bariatric surgery between 2005-2015 in comparison to matched non-surgical controls. In contrast to treatment protocols in tightly controlled randomized clinical trials, real-world treatments, such as those examined by arterburn2020, can evolve over time, especially if comparators include a somewhat loosely defined standard of care (Laiteerapong2015) (e.g., for obesity management), or bariatric surgical procedures where techniques or best practices may change (Sheetz2020, Faria2017, Felsenreich2021, Guldogan2020Sleeve). As such, causal effects of interest in EHR-based studies with long study entry periods may vary over the course of treatment initiation time, which we also refer to as calender time throughout.
Heterogeneity in treatment initiation time has been discussed previously in the context of studies in pharmacoepidemiology. Common across such studies, however, is that variation in treatment initiation time is most often viewed as a potential source of bias towards estimating an implicitly assumed constant treatment effect across calendar time (though for longitudinal or time-to-event outcomes, the effect may vary across time since initiation). For example, stratification of propensity score models by calendar year has been suggested as a solution to address changes in physician prescription patterns or characteristics of treated patients over time (Dusetzina2013, Mack2013, Dilokthornsakul2014CalendarTime). When comparisons of interest involve, say, two drugs approved at different points in time, analysis is often restricted to a period where both medications are available to reduce differences attributable to calendar time (Gagne2013TreatmentDynamics, Suissa2017PrevalentNewUser). Strategies based on instrumental variables (Gokhale2018CalendarTime) and matching (HelinSalmivaara2014SequentialCohort) have also been proposed. In all, the predominant viewpoint has been that variation in causal contrasts by calendar time is something that must be adjusted for, rather than adoption of causal estimands which explicitly use calendar time in their definition. That this is the case likely owes to the fact that in pharmacoepidemiologic studies, typical treatments (e.g., drugs or medications) are not changing over time. In such cases, there may be strong reason to believe there exists a common effect over calendar time in a static population, though the underlying populations may be shifting.
Causal estimands which depend on calendar time have recently been explored in the context of stepped-wedge cluster randomized trials (wang2024model, Lee2025SteppedWedge) and staggered-adoption difference-in-difference settings (callaway2021did, sun2021dynamic), the latter of which often considers analysis of multi-year periods during which a policy is rolled out. Because intervention timing is planned or quasi-experimental across clusters, both study designs provide a natural way to define calendar time-specific effects for groups of clusters initiating treatment at the same time. By contrast, EHR-based studies analyze patient-level outcomes under naturally occurring treatment patterns, which do not follow any controlled schedule, and thus strategies from these literatures do not directly apply.
Determining whether causal effects vary over calender time in EHR-based studies can help inform whether some fundamental aspect of treatment efficacy is changing, and whether researchers might report common or calendar time-varying effect estimates. However, efforts to assess whether treatment efficacy itself is changing are complicated by changing patient populations, and thus, potentially, covariate shift in key effect modifiers. As such, it is equally important to understand why causal effect estimates may vary over time, and the degree to which effect modification plays a role.
In this work, we propose a statistical framework to estimate and describe how and why causal effects vary over calendar time in EHR-based studies. First, doubly-robust, calendar time-specific average treatment effects (ATE) are estimated on the basis of pooled nonparametric nuisance models. Second, estimates are projected onto candidate marginal structural models (MSM), each representing a different pattern in how effect estimates vary across treatment initiation time (e.g., constant, linear etc.). A model selection procedure is then applied to choose which MSM best describes temporal variation in effect estimates. Next, we define a causal estimand that corresponds to the treatment effect that would arise if the distribution of covariates among eligible patients treated at a given point in time instead reflected the distribution observed at some alternative point in time. Finally, we introduce a novel summary metric to quantify the role of covariate shift in explaining observed effect changes and disentangle changes in treatment effects from changes in the patient population receiving treatment. To our knowledge, this work is the first to apply tools from semiparametric theory (tsiatis_semiparametric_2006, kennedy2023semiparametric) to study calendar time-varying treatment effects in EHR-based studies.
This work is motivated by the use of EHR data to study weight loss outcomes following bariatric surgery, particularly in comparison to similarly obese patients who do not undergo surgery, a setting which is described in detail in Section 2, and woven throughout the work. Section 3 formalizes notation and describes calendar time-specific causal estimands. Sections 4 and 5 introduce statistical tools, based on semiparametric efficiency theory, for determining how and why causal effects vary over calendar time, respectively, while Section 6 outlines how analysts can synthesize information from estimation and inferential procedures outlined in previous sections to determine what to report in their own study. Section 7 returns to the motivating study to apply the methods outlined in this work, and finally, Section 8 concludes with discussion. Detailed proofs and simulations which validate tools from this framework are provided in the Supplementary Material.
2 An EHR-based Study of Bariatric Surgery
Bariatric surgery is an intervention to mitigate obesity and related comorbidities, with typical candidates having a body mass index (BMI) exceeding 35 kg/m2 (kaiserpermanente2023). Due to the cost and complexity of conducting large-scale trials for bariatric surgery (Courcoulas2024), EHR-based studies are a critical tool to study long-term outcomes following bariatric surgery, particularly in relation to similarly obese patients not undergoing surgery (arterburn2020, obrien2018microvascular, fisher2018association, Harrington2024, coleman2022bariatric). Despite its success as a weight-loss intervention, only 1-2% of eligible candidates elect to undergo bariatric surgery annually (arterburn2014bariatric). Therefore, EHR-based studies of bariatric surgery must consider a multi-year study entry period to obtain a meaningful number of patients undergoing surgery.
The two most common bariatric surgical procedures are Roux-en-Y gastric bypass (RYGB) and sleeve gastrectomy (SG), with SG having surpassed RYGB in frequency in the past 15 years (arterburn2014bariatric, Reames2014). SG is a less invasive and technically easier procedure than RYGB, and as such is offered by more surgeons. Relevant to this work is that the surgical technique for SG has historically been somewhat less standardized than for RYGB, particularly during the time-frame during which SG became the more commonly performed procedure (Guldogan2020Sleeve, Rosenthal2012, Gagner2013). Given the rapid shifts in the relative popularity of RYGB and SG in recent years, along with possibly evolving surgical techniques and standards of care for obesity management, understanding potential temporal changes in the effectiveness bariatric surgery is of particular interest.
With this backdrop, we examined relative weight change at four pre-specified times post-surgery (6 months, 1 year, 2 years, 3 years), comparing 17,905 patients undergoing bariatric surgery between 2005-2011 at one of three Kaiser Permanente (KP) sites (KP Washington, KP Northern California, KP Southern California) with 933,044 non-surgical patients. In particular, we compared outcomes between surgery and no surgery, each of RYGB/SG versus no surgery separately, and RYGB versus SG. We leveraged EHR data from DURABLE, an NIH-funded study examining long-term outcomes of bariatric surgery across Kaiser Permanente sites (arterburn2020, coleman2016longterm, obrien2018microvascular, coleman2022bariatric, fisher2018association, Harrington2024, haneuse2025).
Using this data source, arterburn2020 previously characterized longitudinal weight trajectories following RYGB and SG in comparison to matched non-surgical patients, though calendar time was considered only as a potential confounder. As an alternative to the matched cohort design employed by arterburn2020, we used a sequential target trial emulation approach (benz2024, madenci2024, haneuse2025, cain2010, danaei2013observational, caniglia2023). Specifically, we considered a sequence of 84 target trials (formalized in Section 3), initiating at monthly intervals between January 2005 and December 2011. Following arterburn2020, our study population included adults aged 19-79 years with BMI 35 kg/m2. We excluded patients without a year of continuous enrollment in KP insurance and those pregnant in the year prior to the start of a particular target trial. 17,596 surgical patients (98.2%) and 763,016 non-surgical patients (81.8%) met the study inclusion criteria at baseline of at least one target trial.
Figure 1: Distribution of patient characteristics in DURABLE electronic health record database over calendar time for eligible patients undergoing bariatric surgery between 2005-2011. Covariate mean values (continuous covariates) or frequency (binary/categorical covariates) are plotted by month of surgery. $m=1$ corresponds to January 2005 and $m=84$ corresponds to December 2011.
On the basis of these data, one might choose to estimate and report common a effect for bariatric surgery on 3 year relative weight change, in comparison to non-surgical patients. Though formal details are deferred until Section 7, such an approach produced estimates of -19.8% (-19.9%, -19.6%) for surgery overall, with corresponding estimates of -23.1% (-23.3%, -22.9%) for RYGB and -16.3% (-16.6%, -16.1%) and SG, in close agreement with estimates of -23.5% and -16.7% reported by arterburn2020, respectively. These summaries, however, do not acknowledge the potential for variation across calendar time, and thus do not offer any insight into possible temporal variation in the effectiveness of bariatric surgery.
Figure 1 summarizes the distribution of several key covariates among eligible bariatric patients across the 84 target trials (i.e., across calendar time). Notably, the distribution of several effect modifiers highlighted by arterburn2020 (age, BMI, race, and diabetes status) varies meaningfully across time of surgery. For example, the mean BMI of patients undergoing bariatric surgery in a given calendar month declines from 47.5 kg/m2 in January 2005 to 43.8 kg/m2 in December 2011, with monthly averages ranging between 42.9 kg/m2 and 49.6 kg/m2. Accordingly, any attempt to study how the effect of bariatric surgery may have changed over the course of the study period should acknowledge and account for this covariate shift. A similar figure displaying covariate shift for non-surgical patients is available in the Supplementary Materials (Figure S1).
3 Problem Setting
3.1 Notation
When a treatment strategy does not correspond to a concrete clinical event, time zero can be challenging to define (hernan2016immortal). In order to ensure that the start of follow-up is well-defined for all patients, particularly those not undergoing surgery, we partition the study entry period into a sequence of $M$ target trials (benz2024, madenci2024, cain2010, danaei2013observational, caniglia2023), indexed by $m\in\{1,\dots,M\}$. In the motivating study of bariatric surgery, $M=84$ and component target trials are conducted at monthly intervals between January 2005 ($m=1$) and December 2011 ($m=M=84$).
To the best of our knowledge, the sequential target trial emulation design has been applied exclusively in applications examining time-to-event outcomes (scola2023implementation). Since relative weight change is not a time-to-event endpoint, we describe notation to formalize this design for continuous outcomes. Let $A_{m}\in\{0,1\}$ denote a binary treatment (e.g., surgery vs. no surgery) received at baseline of trial $m$, $Y_{m}\in\mathbb{R}$ denote an outcome of interest (e.g., relative weight change since baseline of trial $m$), and $\bm{L}_{m}\in\mathbb{R}^{p}$ denote a set of covariates at baseline of trial $m$ assumed sufficient for control of confounding, such as those in Figure 1. Let $E_{m}=g(\bm{L}_{m})$ be a binary eligibility indicator for inclusion in target trial $m$, where $g(\cdot)$ represents a fixed and known eligibility rule (benz2024, benz2025). Practically, because not all patients are in the EHR for the entire duration of the study entry period, we treat subjects as ineligible $(E_{m}=0)$ for periods before they appear and/or after they disappear from the EHR. Thus, at each of $M$ regularly occurring intervals, we ascertain $(\bm{L}_{m},A_{m},Y_{m})$ for eligible subject-trials; formally, we work with coarsened observed data units:
$O=(E_{1},E_{1}\bm{L}_{1},E_{1}A_{1},E_{1}Y_{1},\dots,E_{m},E_{m}\bm{L}_{m},E_{m}A_{m},E_{m}Y_{m},\dots,E_{M},E_{M}\bm{L}_{M},E_{M}A_{M},E_{M}Y_{M}).$
We observe a sample $(O_{1},\ldots,O_{n})$ of $n$ independent and identically distributed observations from an underlying and unknown data distribution, $P$, i.e., the joint distribution of information across both patients and trials. Implicitly throughout, we assume covariates across trials share a common domain $\mathcal{L}$. As will become apparent in Sections 3.2 and 4, covariate, treatment and outcome values of ineligible patients are not required for the definition of our calendar time-specific causal estimands, nor for estimation.
For any $P$-integrable function $f$ we define $Pf=\mathbb{E}_{P}[f]=\int f\,dP$ to denote the expectation of $f$ under $P$, and analogously define $\mathbb{P}_{n}f=\frac{1}{n}\sum_{i=1}^{n}f(O_{i})$ to express the sample mean of $f$ on data units $(O_{1},...,O_{n})$. Lastly, we let $Y_{m}(a_{m})$ denote counterfactual outcomes at time $m$, that is the outcome that would have been observed under treatment $A_{m}=a_{m}$ (hernan2024).
3.2 Calendar Time-Specific Average Treatment Effects
Towards understanding how and why treatment efficacy may vary over treatment initiation time, we define the following causal estimand:
$\chi_{m}(P)=\mathbb{E}[Y_{m}(1)-Y_{m}(0)\penalty 10000\ |\penalty 10000\ E_{m}=1].$
Notice that $\chi_{m}(P)$ is simply the calendar time-specific ATE among patients eligible for trial $m$. Implicit in the definition of $\chi_{m}(P)$ is that comparisons are being made between the specific versions of treatment (i.e., surgical technique, standard of care) in place at time $m$. Note, $\chi_{m}(P)$ is an intention-to-treat (ITT) effect (hernan2016target) in the sense that it captures the effect of initiating treatment at time $m$ $(a_{m}=1)$ versus not initiating $(a_{m}=0)$ without regard to what happens at future time points. In particular, the definition of $\chi_{m}(P)$ is not concerned that treatment definitions may evolve over time (e.g., $a_{j}=0$ and $a_{m}=0$ may mean slightly different things for $j>m$), nor for the possibility of treatment switching.
4 Estimation of Calendar Time-Specific Average Treatment Effects
Given that treatment decisions in EHR-based settings are not randomized, we begin by articulating standard causal inference assumptions to address the potential for confounding. Note that these assumptions are only required to hold among the subset of patients eligible for each trial. This is particularly salient for comparisons of bariatric surgery to no surgery, where study eligibility criteria are often designed to exclude patients who are not realistic candidates for surgery.
A 1 (Consistency)
$Y_{m}(A_{m})=Y_{m}$ when $E_{m}=1$, for all $m$.
A 2 (Positivity)
$\exists\penalty 10000\ \epsilon>0$ such that $\epsilon\leq P(A_{m}=1\penalty 10000\ |\penalty 10000\ \bm{L}_{m},E_{m}=1)\leq 1-\epsilon$, almost surely, for all $m$.
A 3 (No Unmeasured Confounding)
$Y_{m}(a_{m})\perp\!\!\!\perp A_{m}\penalty 10000\ |\penalty 10000\ \bm{L}_{m},E_{m}=1$ for $a_{m}\in\{0,1\}$ and all $m$.
In our data application, we take $\bm{L}_{m}$ to be the most recent values of covariates outlined in Figure 1, although we note that analysts can also include in this vector appropriate summaries of covariate and/or treatment history prior to baseline of trial $m$, if warranted in their application.
4.1 Motivating A Projection Parameter Approach
At the outset, estimation of each calendar time-specific ATE $\chi_{m}(P)$ may seem straightforward. Let $\mu_{m}(a_{m},\bm{L}_{m})=\mathbb{E}[Y_{m}\penalty 10000\ |\penalty 10000\ A_{m}=a_{m},\bm{L}_{m},E_{m}=1]$ and $\pi_{m}(L_{m})=P(A_{m}=1\penalty 10000\ |\penalty 10000\ \bm{L}_{m},E_{m}=1)$ denote the trial $m$ specific outcome regression and propensity score models, respectively. Under Assumptions 1-3, $\chi_{m}(P)$ is identified by the $g$-formula, $\mathbb{E}[\mu_{m}(1,\bm{L}_{m})-\mu_{m}(0,\bm{L}_{m})\penalty 10000\ |\penalty 10000\ E_{m}=1]$ or via inverse probability weighting, $\mathbb{E}\big[\big(\frac{A_{m}}{\pi_{m}(\bm{L}_{m})}-\frac{1-A_{m}}{1-\pi_{m}(\bm{L}_{m})}\big)Y_{m}\penalty 10000\ |\penalty 10000\ E_{m}=1\big]$ (hernan2024). One might then proceed to estimate $\chi_{m}(P)$ on the basis of its nonparametric influence function (tsiatis_semiparametric_2006, kennedy2023semiparametric),
$\dot{\chi}^{*}_{m}(O;P)=\frac{\mathds{1}(E_{m}=1)}{P(E_{m}=1)}$ $\biggr\{\mu_{m}(1,\bm{L}_{m})-\mu_{m}(0,\bm{L}_{m})-\chi_{m}(P)$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ +\biggr(\frac{A_{m}}{\pi_{m}(\bm{L}_{m})}-\frac{1-A_{m}}{1-\pi_{m}(\bm{L}_{m})}\biggr)\Bigr(Y_{m}-\mu_{m}(A_{m},\bm{L}_{m})\Bigr)\biggr\}.$
Unfortunately, estimating $\mu_{m}$ or $\pi_{m}$ by fitting a separate model for each trial would be difficult in practice, especially in settings where a relatively small number of patients initiate a particular treatment during trial $m$. For example, 24 of 84 target trials in our study have fewer than 100 patients undergoing bariatric surgery at the corresponding baseline month, and 12 of 84 of have fewer than 25 surgical patients. This problem is further compounded for comparisons involving the individual surgical procedures, RYGB and SG (Figure S2).
The predominant analytic strategy in sequential target trial emulations for time-to-event outcomes involves fitting a single parametric outcome and/or propensity score model pooling information across trials (scola2023implementation) (e.g., pooled logistic regression (dagostino1990)), perhaps with a spline term to flexibly adjust for calendar time as a confounder (dickerman2019, madenci2024). Given that trends in $\chi_{m}(P)$ over calendar time may reflect a complex interplay between changing treatment efficacy and shifts in effect modifier distributions, applying a pooled parametric model in this setting would require correctly specifying a substantial number of flexible interactions between treatment, confounders, and calendar time. Such a task may be unrealistic in many instances—arguably including our data application given the heterogeneity reported in Figure 1—and might instead motivate the use of pooled models based on flexible machine learning methods.
Ultimately, while we are interested in trial-specific ATEs, we are most interested in the underlying trend in $\chi_{m}(P)$ across calendar time. To formalize estimation of this trend across calendar time, we follow the strategy of kennedy2019 and others (kennedy2023density, neugebauer2007nonparametric, rosenblum2010targeted), and formulate our estimand as a projection of all $M$ trial-specific effects onto a candidate parametric model. To operationalize this approach, we first define a MSM for mean trial-specific counterfactual outcomes.
$\mathbb{E}[Y_{m}(a_{m})\penalty 10000\ |\penalty 10000\ E_{m}=1]=\psi_{0}(m)+\psi(m)a_{m}.$ (1)
When the MSM in Equation (1) is saturated, $\psi(m)=\chi_{m}(P)$. Instead, we adopt a parametric model for $\psi(m)$, denoted $\psi(m;\bm{\beta})$, and take our estimand to be the projection of $\chi_{m}(P)$ onto $\psi(m;\bm{\beta})$. Formally, $\bm{\beta}(P)$ is the projection parameter defined by
$\bm{\beta}(P)=\mathop{\mathrm{argmin}}_{\bm{\beta}\in\mathbb{R}^{k}}\sum_{m=1}^{M}h(m)\Bigr(\chi_{m}(P)-\psi(m;\bm{\beta})\Bigr)^{2},$
where $h(m)$ is an optional trial-specific weight. Henceforth, we proceed with weights of the form $h(m)=w(m)P(E_{m}=1)$ to weight trial contributions according to their effective sample size, where $w(m)$ allows for optional weighting by other additional factors.
This approach offers several advantages, which we elaborate on in subsequent sections. As will become clear in Section 4.2, estimation of $\chi_{m}(P)$ is a necessary step towards estimating $\psi(m;\bm{\beta})$, a task for which we advocate the use of pooled nuisance models based on flexible machine learning methods. Nevertheless, our projection parameter approach allows us to identify meaningful underlying trends in $\chi_{m}(P)$ even if some trial-specific effects are estimated imprecisely (e.g., due to small sample sizes). Although pooled nonparametric models for component nuisance functions will be an essential step in estimating $\chi_{m}(P)$, they provide limited insight into how information is shared across calendar time. Introducing the additional projection step makes this information sharing explicit and clarifies how effect estimates from individual trials are combined to learn trends over calendar time, thereby improving transparency relative to relying on pooled nonparametric models alone. Finally, this approach naturally supports a model selection procedure (described in Section 4.4) for comparing candidate MSMs to identify the functional form that best describes how treatment effects vary over calendar time, including the possibility of a single, constant effect.
4.2 Efficient Estimation of Projection Parameter
We now introduce preliminary efficiency theory for the projection parameters $\bm{\beta}(P)$, which will motivate our proposed estimator. Since influence functions are by definition mean zero, it is helpful to work with uncentered versions, $\dot{\chi}_{m}(O;P)\coloneqq\frac{\mathds{1}(E_{m}=1)}{P(E_{m}=1)}\dot{\chi}^{\dagger}_{m}(O;P)$, which satisfy $\mathbb{E}_{P}[\dot{\chi}_{m}(O;P)]=\chi_{m}(P)$, where
$\dot{\chi}^{\dagger}_{m}(O;P)\coloneqq\mu_{m}(1,\bm{L}_{m})-\mu_{m}(0,\bm{L}_{m})+\biggr(\frac{A_{m}}{\pi_{m}(\bm{L}_{m})}-\frac{1-A_{m}}{1-\pi_{m}(\bm{L}_{m})}\biggr)\Bigr(Y_{m}-\mu_{m}(A_{m},\bm{L}_{m})\Bigr).$
These quantities play an important role in the influence function of $\bm{\beta}(P)$, as summarized in the following theorem.
Theorem 1
Suppose that the candidate MSM, $\psi(m;\bm{\beta})$ is twice differentiable in $\bm{\beta}$. Then, under Assumptions 1-3, $\bm{\beta}(P)$ is identified as the solution to the estimating equation,
$0=\sum_{m=1}^{M}w(m)P(E_{m}=1)\nabla_{\bm{\beta}}\psi(m;\bm{\beta})[\chi_{m}(P)-\psi(m;\bm{\beta})].$
Furthermore, $\dot{\bm{\beta}}^{*}(O;P)$, the influence function of $\bm{\beta}(P)$ (up to proportionality), is given by
$\dot{\bm{\beta}}^{*}(O;P)\propto\dot{\bm{\beta}}^{\dagger}(O;P)=\sum_{m=1}^{M}\mathds{1}(E_{m}=1)w(m)\nabla_{\bm{\beta}}\psi(m;\bm{\beta})|_{\bm{\beta}=\bm{\beta}(P)}\Bigr(\dot{\chi}^{\dagger}_{m}(O;P)-\psi\bigr(m;\bm{\beta}(P)\bigr)\Bigr).$ (2)
[…]
where $\widehat{\chi}_{m}$ is the influence function-based estimator of $\chi_{m}(P)$ mentioned in Section 4.1, and $\widehat{P}$ indicates that any necessary nuisance functions are replaced by estimated quantities. Equations (2) and (3) make clear, however, that to even estimate $\bm{\beta}(P)$, estimates of $\chi_{m}(P)$, and thus estimates of nuisance functions $\mu_{m}$ and $\pi_{m}$, are required. Also of note in Equation (3), trials are weighted proportionally to the sample size of their eligible populations, potentially mitigating the influence of noisier estimates of $\chi_{m}(P)$ in component trials with smaller sample sizes.
[…]
In order to estimate $\chi_{m}(P)$, we first create a pooled dataset $\mathcal{D}$ of patient information across all trials in which they are eligible, that is,
$\mathcal{D}=\bigcup_{m=1}^{M}\Bigr\{(\bm{L}_{m,i},A_{m,i},Y_{m,i},m)\Bigr\}_{i:E_{m,i}=1}.$ (4)
[…]
Routines for estimation of influence function-based estimators like $\widehat{\chi}_{m}$ or $\widehat{\bm{\beta}}$ and typically involve sample splitting or cross-fitting, particularly when using nonparametric and/or machine learning techniques to estimate component nuisance functions (chernozhukov2018). We describe the complete procedure for estimating $\widehat{\chi}_{m}$ and $\widehat{\bm{\beta}}$ below.
Algorithm 1 (Computation of $\widehat{\chi}_{m}$ and $\widehat{\bm{\beta}}$)
Let $S_{1}$ and $S_{2}$ be two disjoint, independent splits of $(O_{1},\ldots,O_{n})$ of sizes $n_{1}$ and $n_{2}$, respectively, with $n=n_{1}+n_{2}$. The analytical procedure to compute the estimator $\widehat{\chi}_{m}$ and $\widehat{\bm{\beta}}$ is as follows:
  1. 1.
    Create pooled dataset of eligible subject-trials $\mathcal{D}_{1}$ among the $n_{1}$ subjects in split $S_{1}$, as in Equation (4).
  2. 2.
[…]
  3. 3.
    For $O_{n_{1}+1},O_{n_{1}+2},...,O_{n}\in S_{2}$, construct plug-in estimates of uncentered influence functions $\{\dot{\chi}_{m}(O_{i};\widehat{P}_{1})\}_{i=n_{1}+1}^{n}$ for each $m\in\{1,\dots,M\}$, where $\dot{\chi}_{m}(O_{i};\widehat{P}_{1})\equiv\frac{\mathds{1}(E_{m}=1)}{\mathbb{P}_{n}[E_{m}]}\dot{\chi}_{m}^{\dagger}(O_{i};\widehat{P}_{1})$.
  4. 4.
    Repeat steps 1-3, swapping the roles of $S_{1}$ and $S_{2}$ to create, for each $m\in\{1,\dots,M\}$, complete sets of influence function contributions from corresponding out-of-sample nuisance function predictions, $\{\dot{\chi}_{m}(O_{i};\widehat{P})\}_{i=1}^{n}=\{\dot{\chi}_{m}(O_{i};\widehat{P}_{1})\}_{i=n_{1}+1}^{n}\cup\{\dot{\chi}_{m}(O_{i};\widehat{P}_{2})\}_{i=1}^{n_{1}}$.
  5. 5.
[…]
  6. 6.
    For a given candidate marginal structural model, $\psi(m;\bm{\beta})$, plug estimates $\{\widehat{\chi}_{m}\}_{m=1}^{M}$ into Equation (3) and solve for $\widehat{\bm{\beta}}$ (e.g., via nleqslv (R-nleqslv)).
4.3 Theoretical Properties of $\widehat{\bm{\beta}}$ and $\widehat{\chi}_{m}$
The choice to estimate $\widehat{\chi}_{m}$ and $\widehat{\bm{\beta}}$ using an estimating equation-based approach rather than a one-step approach is a matter of practical convenience; asymptotically the two approaches are equivalent. In order to describe the asymptotic behavior of $\widehat{\bm{\beta}}$, we note that $\text{Var}[\dot{\bm{\beta}}^{*}(O;P)]=\bm{V}^{-1}\bm{C}\bm{V}^{-1}$, where $\bm{V}\in\mathbb{R}^{k\times k}$ is the scaling matrix of proportionality ignored in Theorem 1, and $\bm{C}=\text{Var}[\dot{\bm{\beta}}^{\dagger}(O;P)]\in\mathbb{R}^{k\times k}$. As we derive in Supplementary Materials Appendix S1, we have that
$\bm{V}=\sum_{m=1}^{M}w(m)P(E_{m}=1)\Bigr\{\nabla_{\bm{\beta}\bm{\beta}}^{2}\psi(m;\bm{\beta})|_{\bm{\beta}=\bm{\beta}(P)}\big[\chi_{m}(P)-\psi\bigr(m;\bm{\beta}(P)\bigr)\big]$ (5)
$-\nabla_{\bm{\beta}}\psi(m;\bm{\beta})|_{\bm{\beta}=\bm{\beta}(P)}[\nabla_{\bm{\beta}}\psi(m;\bm{\beta})|_{\bm{\beta}=\bm{\beta}(P)}]^{T}\Bigr\}.$
Finally, we summarize the asymptotic behavior of $\widehat{\chi}_{m}$ and $\widehat{\bm{\beta}}$ in the following theorem. Henceforth, we write $\lVert f\rVert_{(m)}$ for the $L_{2}(P)$ norm of a function of $(\bm{L}_{m},A_{m},Y_{m})$ conditional on eligibility in the $m^{\text{th}}$ trial: $\lVert f\rVert_{(m)}^{2}\coloneqq\mathbb{E}_{P}[f(\bm{L}_{m},A_{m},Y_{m})^{2}\mid E_{m}=1]$.
Theorem 2
Suppose $\lVert\widehat{\pi}_{m}-\pi_{m}\rVert_{(m)}+\sum_{a=0}^{1}\lVert\widehat{\mu}_{m}(a,\,\cdot\,)-\mu_{m}(a,\,\cdot\,)\rVert_{(m)}=o_{P}(1)$. Further, assume that $\epsilon\leq\widehat{\pi}_{m}(\bm{L}_{m}),\pi(\bm{L}_{m})\leq 1-\epsilon$ and $|Y_{m}|,|\widehat{\mu}_{m}(a,\bm{L}_{m})|\leq\epsilon^{-1}$ almost surely when $E_{m}=1$. Then
$\widehat{\chi}_{m}-\chi_{m}(P)=\mathbb{P}_{n}[\dot{\chi}^{*}_{m}(O;P)]+O_{P}\bigr(\widehat{R}_{m}\bigr)+o_{P}(n^{-1/2}),$
where
$\widehat{R}_{m}=\lVert\widehat{\pi}_{m}-\pi_{m}\rVert_{(m)}\sum_{a=0}^{1}\lVert\widehat{\mu}_{m}(a,\,\cdot\,)-\mu_{m}(a,\,\cdot\,)\rVert_{(m)}.$
Moreover, if $\widehat{R}_{m}=o_{P}(n^{-1/2})$ then $\sqrt{n}\bigr(\widehat{\chi}_{m}-\chi_{m}(P)\bigr)\overset{d}{\to}\mathcal{N}\bigr(0,\textnormal{Var}[\dot{\chi}^{*}_{m}(O,P)]\bigr)$, whereby $\widehat{\chi}_{m}$ attains the nonparametric efficiency bound.
Similarly, under Assumptions S1-S3 in Appendix S1.3,
$\widehat{\bm{\beta}}-\bm{\beta}(P)=\mathbb{P}_{n}[\dot{\bm{\beta}}^{*}(O;P)]+O_{P}\bigg(\sum_{m=1}^{M}\widehat{R}_{m}\bigg)+o_{P}(n^{-1/2}).$
When $\sum_{m=1}^{M}\widehat{R}_{m}=o_{P}(n^{-1/2})$, $\sqrt{n}\bigr(\widehat{\bm{\beta}}-\bm{\beta}(P)\bigr)\overset{d}{\to}\mathcal{N}\bigr(0,\textnormal{Var}[\dot{\bm{\beta}}^{*}(O;P)]\bigr)$, whereby $\widehat{\bm{\beta}}$ attains the nonparametric efficiency bound.
As illustrated by Theorem 2, the rates of convergence for estimates of calendar time-specific treatment effects and their corresponding trend are directly tied to error rates in estimation of the underlying outcome regression and propensity score models. Importantly, convergence rates for both $\widehat{\chi}_{m}$ and $\widehat{\bm{\beta}}$ are faster than component nuisance functions, even when employing flexible modeling choices. When these product errors are $o_{P}(n^{-1/2})$—for example when each nuisance function’s error rate is $o_{P}(n^{-1/4})$—then $\widehat{\chi}_{m}$ is $\sqrt{n}$-consistent and asymptotically normal, and attains the nonparametric efficiency bound. Furthermore, when the required rate conditions are satisfied for each component trial the same properties extend to $\widehat{\bm{\beta}}$. Additional commentary on Assumptions S1-S3 is provided in the Supplementary Materials.
Pooled models can help ensure that the required rate conditions are met in component trials with smaller effective sample sizes. As discussed in Section 4.2, however, pooled parametric models for $\pi_{m}$ and $\mu_{m}$ are likely to be inconsistent in the presence of significant covariate shift, motivating the use of more flexible pooled models. Crucially, when the required rate conditions (e.g., $o_{P}(n^{-1/4})$ for each nuisance function) are satisfied, asymptotically valid inference for the trend in treatment effects across calendar time remains attainable, even when using such flexible pooled models.
In particular, asymptotically valid pointwise Wald-style $(1-\alpha)$-level confidence intervals are given by
[…]
4.4 Model Selection Among Candidate Marginal Structural Models
Suppose one considers a range of candidate MSMs including a constant model (i.e., $\psi(m;\beta)=\beta$), a linear model (i.e., $\psi(m;\bm{\beta})=\beta_{0}+\beta_{1}m$), and more flexible MSMs including higher-order polynomials (in $m$) and splines. A model selection procedure to select the MSM which best describes the underlying trend $\chi_{m}(P)$ is important to best characterize how treatment effects—and thus possibly treatment efficacy—vary over calendar time. In particular, if a constant model were selected, it would suggest that any underlying variation in treatment effects is not detectable given the study sample size, and thus may not be clinically meaningful. On the other hand, if a flexible MSM were selected, it would provide evidence of non-constant trends in treatment effects over calendar time.
To implement such a model selection procedure, we adapt the ideas of kennedy2019 and vanderLaanDudoit2003cv. Let $\{\widehat{\psi}_{1},...,\widehat{\psi}_{K}\}$ denote a set of $K$ candidate MSM projection models, with $\widehat{\psi}_{k}\equiv\psi_{k}(m;\widehat{\bm{\beta}}_{k})$. A natural quantity by which to compare candidate MSMs is mean-squared error (relative to the true $\chi_{m}(P)$). To simplify comparisons, we drop terms that don’t depend on the candidate MSM and proceed with the pseudorisk,
$L(\widehat{\psi}_{k})=\sum_{m=1}^{M}w(m)\bigr\{\psi_{k}(m;\widehat{\bm{\beta}}_{k})^{2}-2\psi_{k}(m;\widehat{\bm{\beta}}_{k})\chi_{m}(P)\bigr\}.$
Given the appearance of $\chi_{m}(P)$ in the pseudorisk, we replace $\chi_{m}(P)$ with estimates based on its influence function, i.e., we proceed with the influence function-based estimator of the pseudorisk,
[…]
This is an intuitive loss function to work with because it reflects an estimating equation-based estimator of the pseudorisk, based on its influence function, treating $\widehat{\psi}_{k}$ as fixed (kennedy2019).
[…]
Figure 2: Overview of cross-fitting procedure for evaluating $\widehat{L}(\widehat{\psi}_{k})$.
In practice, differences between candidate projection models can be small (kennedy2023semiparametric), particularly if models classes are nested within one another. In such cases, we might have a preference for simpler models, such that more flexible MSMs are selected only if there are clinically meaningful trends in $\chi_{m}(P)$ across calendar time. To formalize this, let $\widehat{\psi}^{*}=\mathop{\mathrm{argmin}}_{\widehat{\psi}_{k}\in\{\widehat{\psi}_{1},\dots,\widehat{\psi}_{K}\}}\widehat{L}(\widehat{\psi}_{k})$ denote the MSM among our set of candidates which minimizes the loss function $\widehat{L}(\widehat{\psi}_{k})$. Let
$\epsilon=\sqrt{\frac{\sum_{m=1}^{M}w(m)\mathbb{P}_{n}(E_{m}=1)\widehat{\text{Var}}[\psi^{*}(m,\widehat{\bm{\beta}})]}{\sum_{m=1}^{M}w(m)\mathbb{P}_{n}(E_{m}=1)}}.$
One can interpret $\epsilon$ as a pooled standard deviation across our best estimate of the calendar time trend, weighting proportional to the size of the eligible population at each trial. Finally, one can choose the simplest model such that $|\widehat{L}(\widehat{\psi}_{k})-\widehat{L}(\widehat{\psi}^{*})|\leq c\epsilon$, for some user-selected constant $c$. There is considerable debate in the literature about how large an effect size needs to be on a standard deviation scale to be clinically meaningful (tsujimoto2022mic), though the notion of clinically meaningful is very application dependent. In simulations, we find that $c=1/4$ works well in our setup, but present result for a range of other values. Finally, we note that notion of “simpler model” can be operationalized through any model complexity criterion so long as models can be ordered by this criterion. One reasonable such choice—which we adopt in our application—is ordering candidate MSMs by the dimension of their indexing projection parameter (i.e., the dimension of $\bm{\beta}$).
5 Decomposing Sources of Variation
While the projection parameter framework and corresponding model selection procedure outlined in Section 4 provide insight into how treatment effects vary across calendar time, they fail to explain why such changes (if any) are observed. In particular, if the candidate MSM selected is non-constant, suggesting that $\chi_{m}(P)$ varies over time, we would like to ascertain the degree to which underlying changes in treatment efficacy—as opposed to distribution shift in effect modifiers—are responsible.
5.1 A Cross-Trial Contrast to Remove Distribution Shift
To isolate why treatment effects might vary over calendar time, we begin by defining new contrasts by reweighting mean counterfactual outcomes across eligible populations at different points in time:
$\chi_{j,m}(P)\coloneqq\int_{\mathcal{L}}\mathbb{E}\bigr[Y_{m}(a_{m}=1)-Y_{m}(a_{m}=0)\penalty 10000\ |\penalty 10000\ \bm{L}_{m}=\bm{\ell},g(\bm{L}_{m})=1\bigr]dP_{\bm{L}_{j}|E_{j}=1}(\bm{\ell}\penalty 10000\ |\penalty 10000\ E_{j}=1).$
The estimand $\chi_{j,m}(P)$ represents the difference in mean counterfactual outcomes at time $m$ in a population where trial baseline covariates are standardized to the covariate distribution of the eligible population at time $j$. First, we note that when $j=m$, this contrast reduces to trial-specific ATE defined in Section 3.2, that is, $\chi_{m,m}(P)=\chi_{m}(P)$. Given the definition of the class of estimands $\chi_{j,m}(P)$, one obtains common populations with which to standardize treatment effects across calendar time. Namely, for $m_{1}\neq m_{2}$, differences between $\chi_{j,m_{1}}(P)$ and $\chi_{j,m_{2}}(P)$ cannot be attributable to covariate shift, as the underlying population on which the causal contrast is defined is the same. We exploit this fact in Section 5.3 to introduce a metric for quantifying effect modification.
In order to identify each $\chi_{j,m}(P)$, we introduce an additional assumption similar to those commonly used in transportability analyses (ung2025, dahabreh2019generalizing, dahabreh2020extending).
A 4 (Cross-Trial Overlap)
[…]
Assumption 4 is often referred to as an overlap assumption in the transportability literature and ensures that if a patient met the study eligibility criteria at baseline of trial $j$, their covariate vector $\bm{L}_{j}$ is in the support of the covariate distribution among those eligible for trial $m$.
Lemma 1
Under Assumptions 1-4, $\chi_{j,m}(P)$ is identified by
$\chi_{j,m}(P)=\mathbb{E}[\mu_{m}(1,\bm{L}_{j})-\mu_{m}(0,\bm{L}_{j})\penalty 10000\ |\penalty 10000\ E_{j}=1]=\int_{\mathcal{L}}\{\mu_{m}(1,\bm{\ell})-\mu_{m}(0,\bm{\ell})\}\,dP_{\bm{L}_{j}\mid E_{j}=1}(\bm{\ell}).$
5.2 Estimation of Cross-Trial Effects
Given that the cross-trial effects are just modified versions of calendar time-specific ATEs, the estimation strategy for $\chi_{j,m}(P)$ will be very similar to that of $\chi_{m}(P)$. We define one additional nuisance function, $\xi_{j,m}(\bm{\ell})=\frac{p(\bm{L}_{j}=\ell\penalty 10000\ |\penalty 10000\ E_{j}=1)}{p(\bm{L}_{m}=\ell\penalty 10000\ |\penalty 10000\ E_{m}=1)}$, which represents a density ratio to transport the covariate distribution of the eligible population in trial $m$ to that of trial $j$.
Theorem 3
The influence function of $\chi_{j,m}$ at $P$ in a nonparametric model is given by
$\dot{\chi}^{*}_{j,m}(O;P)=$ $\frac{\mathds{1}(E_{j}=1)}{P(E_{j}=1)}\bigr\{\mu_{m}(1,\bm{L}_{j})-\mu_{m}(0,\bm{L}_{j})-\chi_{j,m}(P)\bigr\}$
$+\frac{\mathds{1}(E_{m}=1)}{P(E_{m}=1)}\xi_{j,m}(\bm{L}_{m})\biggr(\frac{A_{m}}{\pi_{m}(\bm{L}_{m})}-\frac{1-A_{m}}{1-\pi_{m}(\bm{L}_{m})}\biggr)\Bigr(Y_{m}-\mu_{m}(A_{m},\bm{L}_{m})\Bigr).$
The influence function $\dot{\chi}^{*}_{j,m}(O;P)$ admits a very intuitive form. The first summand resembles a $g$-formula estimating function using the outcome regression for trial $m$ on the eligible covariate distribution from trial $j$. The second summand takes the inverse treatment-weighted residuals from eligible patients in trial $m$, and additionally weights them by the appropriate density ratio required to transport to trial $j$. If a subject is eligible for both trials $j$ and $m$, they will contribute in both terms, increasing efficiency, but they can still contribute if eligible for only one of the trials.
Eventually, we will need to estimate $\chi_{j,m}(P)$ for all $M^{2}$ pairs of component target trials. As such, fitting a separate model for each $\xi_{j,m}$ is computationally infeasible. Moreover, conditional density estimation is a challenging statistical problem (wasserman2006all). Because $\xi_{j,m}$ is a ratio of conditional densities, however, we can directly apply Bayes’ rule (sugiyama2012density, benz2025, diaz2023) and leverage pooled models, as with $\widetilde{\mu},\widetilde{\pi}$:
$\xi_{j,m}(\bm{\ell})=\frac{p(\bm{L}_{j}=\bm{\ell}\penalty 10000\ |\penalty 10000\ E_{j}=1)}{p(\bm{L}_{m}=\bm{\ell}\penalty 10000\ |\penalty 10000\ E_{m}=1)}=\frac{P(E_{m}=1)P(T=j\penalty 10000\ |\penalty 10000\ \bm{L}=\bm{\ell},E=1)}{P(E_{j}=1)P(T=m\penalty 10000\ |\penalty 10000\ \bm{L}=\bm{\ell},E=1)},$ (6)
where $T$ denotes the trial membership indicator in full the pooled data set, $\mathcal{D}$. Thus, we have turned the task of conditional density estimation into modeling the trial membership indicator conditional on $\bm{L}$ across all eligible subject-trials. Crucially, this is a classification task well suited for machine learning methods, and we can apply any flexible pooled model $\widetilde{\xi}(m\penalty 10000\ |\penalty 10000\ \bm{\ell})$ to estimate $P(T=m\penalty 10000\ |\penalty 10000\ \bm{L}=\ell,E=1)$.
With additional bookkeeping, we can estimate $\chi_{j,m}(P)$ for all pairs $(j,m)$ within the same routine described by Algorithm 1. First, $\widetilde{\xi}$ is trained during step 2 along with the other pooled models. Following step 3, for each row in $\mathcal{D}$, we save predicted transport ratios to all other trial populations, $\{\widehat{\xi}_{j,m}(\bm{L}_{m})\}_{j=1}^{M}$, and fitted values for the outcome regression at all possible treatment times, $\{(\widehat{\mu}_{m}(1,\bm{L}_{j}),\widehat{\mu}_{m}(0,\bm{L}_{j})\}_{m=1}^{M}$. Extracting $\widehat{\mu}_{m}(a_{m},\bm{L}_{j})=\widetilde{\mu}(a_{m},\bm{L}_{j},m)$ simply requires (artificially) varying the trial index in the pooled model. Upon repeating these steps for both splits of $(O_{1},\dots,O_{n})$, we estimate each cross-trial effect via the following influence function-based estimator:
$\widehat{\chi}_{j,m}$ $=\mathbb{P}_{n}\bigg[\frac{\mathds{1}(E_{j}=1)}{\mathbb{P}_{n}[E_{j}]}\bigr\{\widehat{\mu}_{m}(1,\bm{L}_{j})-\widehat{\mu}_{m}(0,\bm{L}_{j})\bigr\}$
$\quad\quad\quad+\frac{\mathds{1}(E_{m}=1)}{\mathbb{P}_{n}[E_{m}]}\widehat{\xi}_{j,m}(\bm{L}_{m})\biggr(\frac{A_{m}}{\widehat{\pi}_{m}(\bm{L}_{m})}-\frac{1-A_{m}}{1-\widehat{\pi}_{m}(\bm{L}_{m})}\biggr)\Bigr(Y_{m}-\widehat{\mu}_{m}(A_{m},\bm{L}_{m})\Bigr)\bigg],$
where, recall, hatted quantities are estimated in the training split, and $\mathbb{P}_{n}$ represents the empirical mean over the test split.
5.3 Summarizing Changes in Effect Estimates
Ultimately, our interest in $\chi_{j,m}(P)$ is as a stepping stone towards determining the degree to which changes in calendar time-specific treatment effects are explained by changes in effects modifiers compared to something more fundamental in terms of changes in underlying treatment efficacy. We let $\bm{S}\in\mathbb{R}^{M\times M}$ denote the standardization matrix of all $M^{2}$ cross-trial effects, with $S_{jm}=\chi_{j,m}(P)$, and let $\widehat{\bm{S}}$ denote the estimated standardization matrix with entries $\widehat{\chi}_{j,m}$. To quantify sources of variation in cross-trials effects, we define the following variance-like quantities:
$\sigma_{j}^{2}$ $\coloneqq\frac{1}{M-1}\sum_{m=1}^{M}\Bigr(\chi_{j}(P)-\chi_{j,m}(P)\Bigr)^{2},$
$\gamma_{m}^{2}$ $\coloneqq\frac{1}{M-1}\sum_{j=1}^{M}\Bigr(\chi_{m}(P)-\chi_{j,m}(P)\Bigr)^{2}.$
Note that these quantities correspond to the row-wise and column-wise variances of $\bm{S}$, respectively. The quantity $\sigma_{j}^{2}$ characterizes the variability around the trial-specific ATE that would arise if the eligible population at time $j$ had been treated at alternative time points. In contrast, $\gamma_{m}^{2}$ characterizes the variability around $\chi_{m}(P)$ across all trial-eligible populations, had each been treated at time $m$. To quantify the degree to which trends across calendar time might be attributable to covariate shift, we examine the ratio
$\theta=\frac{1}{M}\sum_{m=1}^{M}\theta_{m}=\frac{1}{M}\sum_{m=1}^{M}\frac{\sigma_{m}^{2}}{\sigma_{m}^{2}+\gamma_{m}^{2}}.$
Using our estimated cross-trial effects, $\widehat{\chi}_{j,m}$, we can readily obtain plug-in estimates $\widehat{\sigma}_{m}^{2}$ and $\widehat{\gamma}_{m}^{2}$, and subsequently $\widehat{\theta}$, which offers a useful way to summarize variation across estimated cross-trial effects: $\widehat{\theta}$ can be interpreted as the average proportion of variation in estimated cross-trial effects attributable to factors other than covariate shift, e.g., changes in treatment efficacy. Conditions under which alternatives to changes in treatment efficacy may contribute to variation in cross-trial effects are discussed in greater detail in Section 6.
When $\theta=0$, all $\sigma_{m}^{2}=0$, meaning that within fixed populations, there is no variation in treatment effects across calendar time. In the event that a non-constant candidate MSM is selected, $\theta=0$ indicates that the trend across treatment initiation time is entirely induced by changes in underlying populations. On the other hand, when $\theta=1$, all $\gamma_{m}^{2}=0$, and thus variation in cross-trial effects is fully explained by changing the calendar time index ($m$). Importantly, $\theta=1$ implies that trends across calendar time are not the result of covariate shift. Finally, when $\theta\in(0,1)$, variation in estimated cross-trial effects is partially, but not entirely explained by distribution shift in effect modifiers. Formal hypothesis testing using $\theta$ is discussed in Section 5.5. Complete discussion of using $\theta$, along with $\psi(m;\bm{\beta})$ from Section 4, to explain both how and why causal effects may vary over calendar time, is presented in Section 6.
5.4 Theoretical Properties of $\widehat{\chi}_{j,m}$
We briefly summarize the asymptotic behavior of $\widehat{\chi}_{j,m}$, which will be necessary for inference on $\theta$.
Theorem 4
Suppose $\lVert\widehat{\xi}_{j,m}-\xi_{j,m}\rVert_{(m)}+\lVert\widehat{\pi}_{m}-\pi_{m}\rVert_{(m)}+\sum_{a=0}^{1}\lVert\widehat{\mu}_{m}(a,\,\cdot\,)-\mu_{m}(a,\,\cdot\,)\rVert_{(m)}=o_{P}(1)$. Further, assume that $\epsilon\leq\widehat{\pi}_{m}(\bm{L}_{m}),\pi(\bm{L}_{m})\leq 1-\epsilon$ and $|Y_{m}|,|\widehat{\mu}_{m}(a,\bm{L}_{m})|,|\xi_{j,m}|,|\widehat{\xi}_{j,m}|\leq\epsilon^{-1}$ almost surely when $E_{m}=1$. Then
$\widehat{\chi}_{j,m}-\chi_{j,m}(P)=\mathbb{P}_{n}[\dot{\chi}^{*}_{j,m}(O;P)]+O_{P}\bigr(\widehat{R}_{j,m}\bigr)+o_{P}(n^{-1/2}),$
where
$\widehat{R}_{j,m}=\sum_{a=0}^{1}\lVert\widehat{\mu}_{m}(a,\,\cdot\,)-\mu_{m}(a,\,\cdot\,)\rVert_{(m)}\left\{\lVert\widehat{\xi}_{j,m}-\xi_{j,m}\rVert_{(m)}+\lVert\widehat{\pi}_{m}-\pi_{m}\rVert_{(m)}\right\}.$
Moreover, if $\widehat{R}_{j,m}=o_{P}(n^{-1/2})$ then $\sqrt{n}\bigr(\widehat{\chi}_{j,m}-\chi_{j,m}(P)\bigr)\overset{d}{\to}\mathcal{N}\bigr(0,\textnormal{Var}[\dot{\chi}^{*}_{j,m}(O,P)]\bigr)$, whereby $\widehat{\chi}_{j,m}$ attains the nonparametric efficiency bound.
The asymptotic behavior of $\widehat{\chi}_{j,m}$ is very similar to that of $\widehat{\chi}_{m}$, with an additional requirement that the product error between transport model $\xi_{j,m}$ and outcome model $\mu_{m}$ is $o_{P}(n^{-1/2})$. This rate condition is not unreasonable in light of Equation (6), and the choice to model trial index probabilities rather than conditional densities directly. We will leverage the asymptotic normality of $\widehat{\chi}_{j,m}$ in the following section when conducting inference for $\theta$.
5.5 Hypothesis Testing for $\theta$
Consider the hypothesis test $H_{0}:\theta\in\{0,1\}$ versus $H_{1}:\theta\in(0,1)$. We frame the null hypothesis as simultaneously testing two point nulls because $\theta=0$ and $\theta=1$ correspond to the cases where all or none of the variation in cross-trial effects, respectively, is explained by changes in underlying populations, respectively. Rejection of $H_{0}$ for this test is equivalent to simultaneous rejection of the two one-sided tests, $H_{0}^{(0)}:\theta=0$ versus $H_{1}^{(0)}:\theta>0$ and $H_{0}^{(1)}:\theta=1$ versus $H_{1}^{(1)}:\theta<1$. Drivers of treatment effect heterogeneity across calendar time differ based on whether $\theta=1$ or $\theta\in(0,1)$, and thus only testing the null $H_{0}^{(0)}$ is insufficient. Moreover, as discussed in Section 6, rejection of $H_{0}^{(1)}$ in favor of $\theta<1$ may motivate different standardization strategies when the primary goal is to understand changes in treatment efficacy.
In practice, $H_{0}:\theta\in\{0,1\}$ considers testing at both boundary points for $\theta$, somewhat similar to testing variance components in mixed-effects models (stram1994variance) and as such characterizing the asymptotic distribution of $\widehat{\theta}$ under the null would likely yield a challenging ratio involving mixtures of chi-squared distributions. On the other hand, a full nonparametric bootstrap is computationally infeasible in our setting, given that the size of the pooled dataset $\mathcal{D}$ exceeds several million eligible subject-trials for comparisons involving non-surgical patients (Figure S2). Not only does re-estimation of nuisance functions across bootstrapped datasets pose a large computational burden, but so too does prediction of $\widehat{\mu}_{m}(a_{m},\bm{L}_{j})$ and $\widehat{\xi}_{j,m}(\bm{L}_{m})$ given that the number of predictions required is a $M$-fold larger than the size of $\mathcal{D}$.
Instead, we use an asymptotic version of the parametric bootstrap (van2000asymptotic, EfronTibshirani1994) for $\bm{S}$. In particular, under the conditions of Theorem 4, $\sqrt{n}\bigr(\widehat{\bm{S}}-\bm{S})\overset{d}{\to}\mathcal{N}(0,\bm{\Sigma})$ where $\bm{\Sigma}\in\mathbb{R}^{M^{2}\times M^{2}}$ is the covariance matrix of $\bm{S}$, with entries $\bm{\Sigma}(s_{m_{1}j_{1}},s_{m_{2}j_{2}})=\text{Cov}[\dot{\chi}^{*}_{j_{1},m_{1}}(O;P),\dot{\chi}^{*}_{j_{2},m_{2}}(O;P)]$. Critically, this asymptotic normality for $\bm{S}$ holds even when $\theta$ is at a boundary point. Letting $\widehat{\bm{\Sigma}}_{n}$ denote the empirical covariance matrix of cross-trial influence function contributions, we resample $\widehat{\bm{S}}^{(1)},\dots,\widehat{\bm{S}}^{(B)}\sim\mathcal{N}(\widehat{\bm{S}},\widehat{\bm{\Sigma}}_{n})$. Finally, to approximate the sampling distribution of $\widehat{\theta}$, summarize resamples $\widehat{\bm{S}}^{(1)},\dots,\widehat{\bm{S}}^{(B)}$ by their corresponding variance ratios $\widehat{\theta}^{(1)},\dots,\widehat{\theta}^{(B)}$.
Unfortunately, due to the boundary null, bootstrapped replicates $\widehat{\theta}^{(1)},\dots,\widehat{\theta}^{(B)}$ will never be exactly 0 or 1 and thus 95% percentile-based confidence intervals would never reject $H_{0}$. We follow a similar strategy to that of CavaliereNielsenPedersenRahbek2022, and modify our test to allow for a small margin away from the boundary. For some small $\delta>0$, we adopt the test $H_{0}^{(\delta)}:\theta\in[0,\delta]\cup[1-\delta,1]$ versus $H_{1}^{(\delta)}:\theta\in(\delta,1-\delta)$, and reject $H_{0}^{(\delta)}$ if a 95% percentile-based confidence interval is fully within $(\delta,1-\delta)$. One can interpret $\delta$ as the fraction of variability in entries in $\bm{S}$ for which relative changes in causal effects is deemed negligible. The choice of $\delta$ is highly context dependent, as what constitutes clinically meaningful changes is application specific; in the present context of bariatric surgery, we adopt $\delta=0.05$.
6 Practical Guidance for How and Why Effects Vary
Thus far, we have introduced statistical tools to characterize how treatment effects change across the study entry period, along with summary measures to quantify the role of covariate shift in explaining why effects might change. In our experience, each set of tools in isolation reflects an incomplete picture, and both are necessary to inform analysts how best to understand and report findings. Although the goals of each study differ, and we do not advocate for a “one size fits all” approach, Section 6.1 and Table 1 summarize how analysts might use tools from this work to help interpret, characterize, and report calendar time-varying effects in their study.
Under Assumptions 1-4 and the conditions of Theorems 2 and 4, the only possible explanations for changes in treatment effects over calendar time are due to covariate shift in effect modifiers or changes in underlying treatment efficacy. Section 6.2 comments on alternative explanations under violations of these assumptions.
6.1 Guidance Under No Violations of Assumptions 1-4
To illustrate how model selection and hypothesis testing for $\theta$ can be used together to understand reporting of calendar time-varying effects, we consider two scenarios, one where a constant MSM is selected and a second where a non-constant MSM is selected. First, when the selected MSM is a constant, it is reasonable to report a common effect for the study population, $\psi(m;\widehat{\bm{\beta}})=\widehat{\beta}$, regardless of the value of $\theta$. In some settings, $H_{0}^{(0)}:\theta=0$ might be rejected despite a constant MSM having been selected. Such cases typically arise when differences in estimated treatment effects across calendar time are very small and manifest as random deviations centered around a common effect. Altogether, selecting a constant MSM indicates minimal heterogeneity in the underlying effect over calendar time, suggesting that any observed variation is unlikely to be clinically meaningful.
[…]
Projection Model: Constant $\bigr(\psi(m,\bm{\beta})=\beta\bigr)$
Fail to Reject $H_{0}^{(\delta)}$ ($\theta\leq\delta$) Fail to Reject $H_{0}^{(\delta)}$ ($\theta\geq 1-\delta$) Reject $H_{0}^{(\delta)}$($\theta\in(\penalty 10000\ \delta,1-\delta\penalty 10000\ )$)
Calendar Time Varying Effect (Study Population)
Calendar Time Varying Effect (Fixed Population) $\checkmark^{*}$ $\checkmark^{*}$
Reason(s) for Variation No Variation Variation not clinically meaningful Variation not clinically meaningful
Action Report common effect Report common effect Report common effect
Projection Model: Not Constant $\bigr(\psi(m,\bm{\beta})\neq\beta\bigr)$
Fail to Reject $H_{0}^{(\delta)}$ ($\theta\leq\delta$) Fail to Reject $H_{0}^{(\delta)}$ ($\theta\geq 1-\delta$) Reject $H_{0}^{(\delta)}$ ($\theta\in(\delta,1-\delta)$)
Calendar Time Varying Effect (Study Population)
Calendar Time Varying Effect (Fixed Population) $\checkmark$ $\checkmark$
Reason(s) for Variation Covariate shift in effect modifiers Possible changes in treatment efficacy Covariate shift in effect modifiers, possible changes in treatment efficacy
Action Acknowledge changes across time driven by changes in underlying populations
Consider standardization to fixed population and reporting common effect in that population
Report calendar time-varying effect Report calendar time-varying effect
Consider standardization to fixed population to further study changes in treatment efficacy
Table 1: Summary of possible reporting decisions based on joint results from projection marginal structural model and $\theta$.
6.2 Guidance under Unmeasured Confounding or Model Misspecification
Including $m$ as a covariate in nonparametric pooled models can make it difficult to disentangle differential forms of bias across trials from changes in the treatment effect over calendar time, i.e., it can be that $\widehat{\theta}$ is greater than 0 even if the underlying treatment efficacy itself is not changing. Under violations of Assumptions 3, for example, the difference between $\widehat{\chi}_{j,m_{1}}$ and $\widehat{\chi}_{j,m_{2}}$ would reflect both the difference in treatment efficacy between times $m_{1}$ and $m_{2}$ but also the difference in bias introduced by unmeasured confounding. A similar phenomenon holds under nuisance model misspecification such that trial index covariate captures residual confounding. It is exactly for these reasons why reliance on $\theta$ on its own is inadequate to characterize changing treatment efficacy. Though biases may introduce small differences between $\widehat{\chi}_{j,m_{1}}$ and $\widehat{\chi}_{j,m_{2}}$ even when $\chi_{j,m_{1}}(P)=\chi_{j,m_{2}}(P)$ (i.e. producing $\widehat{\theta}>0$), a constant projection MSM will still be selected if differences between treatment effect estimates across calendar time are insubstantial. Ultimately if there is a strong concern of wrongful attribution of credit towards changes in underlying treatment efficacy, one can choose a larger value of $c$ in the projection MSM selection procedure, thereby enforcing that changes in $\widehat{\chi}_{m}$ over time must be greater in order for a non-constant MSM to be selected.
6.3 Simulations
Extensive simulations, tied closely to the motivating study of bariatric surgery, which validate the utility of the tools outlined in Sections 4 and 5, are provided in Appendix S3. Simulations highlight how MSM model selection varies by sample size and choice of shrinkage parameter $c$, along with the ability of the inferential procedure described in 5.5 to distinguish between effect modification from underlying variation in treatment efficacy.
7 Application of Framework to Bariatric Surgery
Finally, we return to our motivating study of bariatric surgery and apply the proposed methods to examine trends over calendar time from 2005 to 2011. Analyses were conducted across 16 settings, defined by four post-surgical follow-up times (6 months, 1 year, 2 years, and 3 years) and four treatment comparisons (no surgery versus surgery, no surgery versus RYGB or SG separately, and RYGB versus SG). No SG patients met study eligibility criteria until month $m=30$, so analysis for comparisons involving SG was restricted to trials $m\in\{30,\dots 84\}$.
Random forests were used to estimate all pooled nuisance models, separately for each setting, adjusting for age, BMI, sex, race, KP study site, self-reported smoking status, hypertension, dyslipdeimia, type II diabetes, and prescriptions for antilipemic or antihypertensive medications, or insulin. This set of potential confounders directly followed from those used by arterburn2020. Candidate MSMs for projection included constant, linear, and cubic models, and splines with 2 and 3 internal knots. To select among candidate MSMs, we employed the procedure described in Section 4.4 with $c=0.25$. Additional details on precise covariate and outcome definitions, as well as results for other choices of $c$, are provided in the Supplementary Materials.
[…]
Based on selected MSMs, surgical patients experienced excess 3-year weight change (relative to non-surgical patients) of -27.4% (-30.1%, -24.8%) at $m=1$ and -18.3% (-18.8%, -17.8%) at $m=84$, reflecting a change of 9.1% estimated treatment effect between 2005 and 2011. This temporal shift indicates that the common effect of –19.6% reported in Section 2, obtained by projection onto a constant MSM, obscures an important aspect understanding the efficacy of bariatric surgery.
Similar changes were observed for RYGB patients, with estimated 3-year weight change differences ranging from -28.3% (-31.3%, -25.3%) at $m=1$ to -23.0% (-23.5%, -22.5%) at $m=84$. As reflected in Figure 3, there is little change in the 3-year RYGB effect estimate after $m=20$. Unsurprisingly then, the common effect estimate of -23.1% reported in Section 2 is comprises a reasonable summary of the treatment effect for a much greater portion of the study period.
A constant MSM was selected for the 3-year weight change outcome comparing SG patients relative to non-surgical patients. Thus, the reported estimate of -16.3% from Section 2 already reflects what one might reasonably report, after considering the possibility of variation over calendar time. Non-constant MSMs were selected in other comparisons involving SG patients, however, as 6-month weight change treatment effects for SG patients between $m=30$ and $m=84$ improved by 1.0% [-17.5% (-17.7%, -17.2%) to -18.5% (-19.0%, -17.9%)] compared to non-surgical patients, and by 1.3% [4.7 (4.0, 5.5)% to 3.4% (2.8%, 4.0%)] compared to patients undergoing RYGB.
Figure 3B presents estimates of $\widehat{\chi}_{j,m}$ for select comparisons. For a fixed $m$, variation in the vertical direction reflects changes in cross-trial effects when varying the underlying population (captured by $\widehat{\gamma}_{m}^{2}$). Conversely, for a fixed color, variation in the height of points reflects changes in cross-trial effects in a particular population across calendar time (captured by $\widehat{\sigma}_{m}^{2}$). Notably, the greater degree of variation in the vertical direction at fixed points in time (i.e., when varying populations) in the top row of Figure 3B suggests that covariate shift plays a larger role in explaining trends across calendar time for comparisons of no surgery to all surgical procedures than to SG alone. This is further reflected by 95% confidence intervals for $\theta$ in Table 2, which reject $H_{0}^{(\delta)}$ in all comparisons except those between SG and no surgery (with $\delta=0.05$).
Comparison Outcome Minimizing MSM Selected MSM 95% Interval for $\theta$
6 Months Spline (3 Knots) Cubic (0.684, 0.809)
1 Year Spline (3 Knots) Spline (3 Knots) (0.648, 0.767)
2 Years Spline (2 Knots) Spline (2 Knots) (0.644, 0.765)
Surgery vs. No Surgery 3 Years Spline (3 Knots) Spline (3 Knots) (0.682, 0.777)
6 Months Cubic Cubic (0.668, 0.821)
1 Year Spline (3 Knots) Spline (3 Knots) (0.658, 0.795)
2 Years Spline (3 Knots) Spline (3 Knots) (0.647, 0.785)
RYGB vs. No Surgery 3 Years Spline (3 Knots) Spline (3 Knots) (0.670, 0.787)
6 Months Spline (3 Knots) Spline (3 Knots) (0.971, 0.989)
1 Year Constant Constant (0.974, 0.991)
2 Years Linear Constant (0.980, 0.992)
SG vs. No Surgery 3 Years Constant Constant (0.979, 0.992)
6 Months Spline (3 Knots) Cubic (0.718, 0.807)
1 Year Constant Constant (0.670, 0.784)
2 Years Constant Constant (0.666, 0.784)
RYGB vs. SG 3 Years Constant Constant (0.688, 0.793)
Table 2: Summary of model selection results and 95% bootstrapped intervals for $\theta$. The minimizing MSM denotes $\widehat{\psi}^{*}$ minimizing $\widehat{L}(\widehat{\psi}$), while the selected MSM denotes the simplest model within $c=0.25$ weighted standard deviations of the minimizer.
It is not overly surprising that underlying treatment efficacy for bariatric surgery relative to no surgery appears to have “declined” over time, given that the composition of bariatric procedures, specifically the proportion of RYGB and SG, has itself changed over time. Nevertheless, that our method is able to discover this well-explained trend offers evidence of its utility. When making comparisons to non-surgical patients, much of the clinical literature aggregates all surgical procedures under a single label “bariatric surgery” (obrien2018microvascular, Harrington2024, benz2024, madenci2024, haneuse2025, fisher2018association). As such, recognizing that both the meaning and the underlying effectiveness of this aggregated treatment label have changed is particularly relevant. Our results also hint at possible improvements in SG technique and improved weight loss outcomes 6-months post surgery, though such improvements are not observed for longer-term outcomes. One possible explanation is that any underlying changes in technique have helped patients reach their weight-loss nadir more quickly, but have not influenced peak weight-loss or rebound effects (courcoulas2013weight, arterburn2020).
8 Discussion
Observational studies leveraging EHR data are inherently subject to many factors beyond the control of investigators, particularly with respect to how treatments are administered and to whom. When a study reports a single causal effect (e.g., “the” ATE), that estimate implicitly reflects an average of calendar time-specific effects, which may not themselves be constant over time. Although the appropriate structure of reported effects depends largely on the goals of a given study, the methods in this work are designed to clarify both how and why causal effects may vary across calendar time. In doing so, the intent is that they will help guide decisions about what to report, particularly when potential changes to underlying treatment efficacy over time are of interest.
One key benefit of the sequential target trial approach considered in this work is that the design directly establishes initiator cohorts (i.e., each with a common time zero), through which calendar time-specific treatment effects can be defined. The proposed methods generalize to other study designs, however, provided that the study entry period can be discretized into sufficiently narrow windows through which one can define calendar time-specific effects. For example, head-to-head comparisons of RYGB and SG do not strictly require a sequence of target trials because the start of follow-up is defined by the date of surgery, and a (matched) cohort design may be a reasonable alternative (McTigue2020, benz2025). Under such designs, one might partition the study entry period into small intervals and define contrasts analogous to $\chi_{m}(P)$ and $\chi_{j,m}(P)$, before applying the ideas introduced in this work. The choice of interval length will be application specific, and depend on both how quickly clinical practice can change as well as the frequency of treatment.
In studies where there is no clinical rationale for changes in underlying treatment efficacy (e.g., when treatments are fixed-dose pharmaceutical medications), summarizing variation in cross-trial effects can serve as a useful diagnostic for detecting effect modification over time. For instance, $\theta$ may be particularly informative in pharmacoepidemiologic studies, helping to quantify how treatment effects vary as the characteristics of medication users evolve with broader adoption.
We note that there are natural connections between our work and prior work on violations of the consistency assumption in causal inference (cole2009consistency, vanderweele2009consistency, VanderWeele2013). While RYGB and SG can be framed as multiple versions of bariatric surgery, it is perhaps less natural to view each procedure over time as multiple versions of treatment in the usual sense, in part because it is rarely known when treatments differ, and thus which “version” of treatment was received. In a sense, our setup implicitly defines trial-specific versions of treatment, i.e., through the symbols $A_{m}$ and $a_{m}$. However, we choose to view counterfactual outcomes $Y_{m}(a_{m})$ similarly to those defined and framed in cole2009consistency, $Y(a,k)$, where the additional treatment details $(k)$ required to maintain consistency are entirely captured by trial index ($m$) in our notation. In particular, we adopt the viewpoint that if $a_{m}=a_{j}$ but $Y_{m}(a_{m})\neq Y_{j}(a_{j})$, then this reflects changes in the efficacy of a particularly treatment $a$ between two time points $m,j$, rather than always calling this phenomenon multiple versions of treatment. This choice is in line with Cole and Frangakis, who note that taking versions of treatment to the extreme precludes any sort of generalizability and hampers the utility of a study.
Finally, we have assumed throughout that during the time period where a patient is enrolled in Kaiser Permanente insurance, and thus is accessible in the EHR, there is no missing data in components of $O$. In practice, missing covariates may preclude ascertainment of eligibility, or outcomes may be missing even when eligibility is ascertainable. For the purposes of this work, we treat patients with incomplete information at a particular trial as ineligible $(E_{m}=0)$. However, we see combining ideas from our framework with recent work on causal inference under missing eligibility (benz2024, benz2025) as key area of future research. Simultaneously handling possibly missing treatment status and/or outcomes across trials represents another important line of further methodological development.
Code Availability
All code for analysis and simulations is made available on GitHub at https://github.com/lbenz730/calendar_time.
Acknowledgments
This work was supported by NIH Grants R01 DK128150-01, R01 HL166324, and F31 DK141237-01

### v2

@@
 Abstract
-Comparative effectiveness studies using electronic health records (EHR) consider data from patients who could “enter” the study cohort at any point during an interval that spans many years in calendar time.
-Unlike treatments in tightly controlled trials, real-world treatments can evolve over calendar time, especially if comparators include standard of care, or procedures where techniques may improve.
-Efforts to assess whether treatment efficacy itself is changing are complicated by changing patient populations, with potential covariate shift in key effect modifiers.
-In this work, we propose a statistical framework to estimate calendar-time specific average treatment effects and describe both how and why effects vary across treatment initiation time in EHR-based studies.
+Standard practice in electronic health record (EHR)-based studies evaluating the comparative effectiveness of bariatric surgery relative to no surgery is to estimate and report a constant treatment effect across calendar time.
+However, real-world treatment strategies can evolve, particularly when comparators include standard of care or surgical procedures where techniques may improve, making it clinically important to ascertain whether efficacy of bariatric surgery has changed over time.
+Efforts to determine whether treatment efficacy itself is evolving are complicated by changing patient populations, with potential covariate shift in key effect modifiers.
+Through a comprehensive analysis of EHR data from Kaiser Permanente following two bariatric surgical procedures compared to standard of care, we develop a statistical framework to estimate calendar time-specific average treatment effects and describe both how and why effects vary across treatment initiation time in EHR-based studies.
 Our approach projects doubly robust, time-specific treatment effect estimates onto candidate marginal structural models and uses a model selection procedure to best describe how effects vary by treatment initiation time.
 We further introduce a novel summary metric, based on standardization analysis, to quantify the role of covariate shift in explaining observed effect changes and disentangle changes in treatment effects from changes in the patient population receiving treatment.
-Extensive simulations using EHR data from Kaiser Permanente are used to validate the utility of the framework, which we apply to study changes in relative weight loss following two bariatric surgical interventions versus no surgery among patients with severe obesity between 2005–2011.
+Keywords: causal inference, treatment effect heterogeneity, bariatric surgery, electronic health records
 1 Introduction
-Electronic health record (EHR) databases provide a useful data source for comparative effectiveness studies, allowing for larger sample sizes from patient populations which are more diverse than those typical of clinical trials (Casey2016EHR).
-In EHR-based studies, study entry periods frequently extend over a decade or more (arterburn2020, fisher2018association, coleman2016longterm, Harrington2024, haneuse2025, o2013long), perhaps because (a) treatment initiation is uncommon, (b) long-term outcomes are of interest, and/or (c) extending data collection further back in time requires little additional effort.
-For example, in a large EHR-based study, arterburn2020 examined long-term weight loss outcomes of patients undergoing bariatric surgery between 2005-2015 in comparison to matched non-surgical controls.
-In contrast to treatment protocols in tightly controlled randomized clinical trials, real-world treatments, such as those examined by arterburn2020, can evolve over time, especially if comparators include a somewhat loosely defined standard of care (Laiteerapong2015) (e.g., for obesity management), or bariatric surgical procedures where techniques or best practices may change (Sheetz2020, Faria2017, Felsenreich2021, Guldogan2020Sleeve).
-As such, causal effects of interest in EHR-based studies with long study entry periods may vary over the course of treatment initiation time, which we also refer to as calender time throughout.
+Electronic health record (EHR) databases provide a useful data source for comparative effectiveness studies, allowing for larger sample sizes from patient populations which are more diverse than those typical of clinical trials.
+In EHR-based studies, study entry periods frequently extend over a decade or more (Arterburn et al., 2020, Fisher et al., 2018, Coleman et al., 2016, Harrington et al., 2024, Haneuse et al., 2025), perhaps because (a) treatment initiation is uncommon, (b) long-term outcomes are of interest, and/or (c) extending data collection further back in time requires little additional effort.
+For example, in a large EHR-based study, Arterburn et al.
+Arterburn et al.
+(2020) examined long-term weight loss outcomes of patients undergoing bariatric surgery between 2005-2015 in comparison to matched non-surgical controls.
+In contrast to treatment protocols in tightly controlled randomized clinical trials, real-world treatments, such as those examined by Arterburn et al.
+Arterburn et al.
+(2020), can evolve over time, especially if comparators include a somewhat loosely defined standard of care (Laiteerapong and Huang, 2015) (e.g., for obesity management), or bariatric surgical procedures where techniques or best practices may change (Sheetz et al., 2020, Faria, 2017, Felsenreich et al., 2021, Guldogan et al., 2020).
+As such, causal effects of interest in EHR-based studies with long study entry periods may vary over the course of treatment initiation time, which we also refer to as “calendar time” throughout.
 Heterogeneity in treatment initiation time has been discussed previously in the context of studies in pharmacoepidemiology.
 Common across such studies, however, is that variation in treatment initiation time is most often viewed as a potential source of bias towards estimating an implicitly assumed constant treatment effect across calendar time (though for longitudinal or time-to-event outcomes, the effect may vary across time since initiation).
-For example, stratification of propensity score models by calendar year has been suggested as a solution to address changes in physician prescription patterns or characteristics of treated patients over time (Dusetzina2013, Mack2013, Dilokthornsakul2014CalendarTime).
-When comparisons of interest involve, say, two drugs approved at different points in time, analysis is often restricted to a period where both medications are available to reduce differences attributable to calendar time (Gagne2013TreatmentDynamics, Suissa2017PrevalentNewUser).
-Strategies based on instrumental variables (Gokhale2018CalendarTime) and matching (HelinSalmivaara2014SequentialCohort) have also been proposed.
+For example, stratification of propensity score models by calendar year has been suggested as a solution to address changes in physician prescription patterns or characteristics of treated patients over time (Dusetzina et al., 2013, Mack et al., 2013, Dilokthornsakul et al., 2014).
+When comparisons of interest involve, say, two drugs approved at different points in time, analysis is often restricted to a period where both medications are available to reduce differences attributable to calendar time (Gagne et al., 2013, Suissa et al., 2017).
+Strategies based on instrumental variables (Gokhale et al., 2018) and matching (Helin-Salmivaara et al., 2014) have also been proposed.
 In all, the predominant viewpoint has been that variation in causal contrasts by calendar time is something that must be adjusted for, rather than adoption of causal estimands which explicitly use calendar time in their definition.
@@
 In such cases, there may be strong reason to believe there exists a common effect over calendar time in a static population, though the underlying populations may be shifting.
-Causal estimands which depend on calendar time have recently been explored in the context of stepped-wedge cluster randomized trials (wang2024model, Lee2025SteppedWedge) and staggered-adoption difference-in-difference settings (callaway2021did, sun2021dynamic), the latter of which often considers analysis of multi-year periods during which a policy is rolled out.
+Causal estimands which depend on calendar time have recently been explored in the context of stepped-wedge cluster randomized trials (Wang et al., 2024, Lee et al., 2025) and staggered-adoption difference-in-difference settings (Callaway and Sant’Anna, 2021, Sun and Abraham, 2021), the latter of which often considers analysis of multi-year periods during which a policy is rolled out.
 Because intervention timing is planned or quasi-experimental across clusters, both study designs provide a natural way to define calendar time-specific effects for groups of clusters initiating treatment at the same time.
 By contrast, EHR-based studies analyze patient-level outcomes under naturally occurring treatment patterns, which do not follow any controlled schedule, and thus strategies from these literatures do not directly apply.
-Determining whether causal effects vary over calender time in EHR-based studies can help inform whether some fundamental aspect of treatment efficacy is changing, and whether researchers might report common or calendar time-varying effect estimates.
-However, efforts to assess whether treatment efficacy itself is changing are complicated by changing patient populations, and thus, potentially, covariate shift in key effect modifiers.
+Determining whether causal effects vary over calendar time in EHR-based studies is of clinical interest as such efforts can help characterize the real-world history of a treatment and inform whether some fundamental aspect of its efficacy is changing.
+Moreover, describing the structure of causal effects over calendar time yields an estimate of the treatment effect at the end of the study period (i.e., that which most closely reflects current practice) and informs how historical data might be leveraged to guide present-day clinical decision-making.
+However, such efforts are complicated by changing patient populations, and thus, potentially, covariate shift in key effect modifiers.
 As such, it is equally important to understand why causal effect estimates may vary over time, and the degree to which effect modification plays a role.
-In this work, we propose a statistical framework to estimate and describe how and why causal effects vary over calendar time in EHR-based studies.
+In this work, motivated by our experience in a series of studies of long-term outcomes following bariatric surgery (Arterburn et al., 2020, Coleman et al., 2016, O’Brien et al., 2018, Coleman et al., 2022, Fisher et al., 2018, Harrington et al., 2024, Haneuse et al., 2025), we propose a statistical framework to estimate and describe how and why causal effects vary over calendar time in EHR-based studies.
 First, doubly-robust, calendar time-specific average treatment effects (ATE) are estimated on the basis of pooled nonparametric nuisance models.
-Second, estimates are projected onto candidate marginal structural models (MSM), each representing a different pattern in how effect estimates vary across treatment initiation time (e.g., constant, linear etc.).
-A model selection procedure is then applied to choose which MSM best describes temporal variation in effect estimates.
-Next, we define a causal estimand that corresponds to the treatment effect that would arise if the distribution of covariates among eligible patients treated at a given point in time instead reflected the distribution observed at some alternative point in time.
-Finally, we introduce a novel summary metric to quantify the role of covariate shift in explaining observed effect changes and disentangle changes in treatment effects from changes in the patient population receiving treatment.
-To our knowledge, this work is the first to apply tools from semiparametric theory (tsiatis_semiparametric_2006, kennedy2023semiparametric) to study calendar time-varying treatment effects in EHR-based studies.
-This work is motivated by the use of EHR data to study weight loss outcomes following bariatric surgery, particularly in comparison to similarly obese patients who do not undergo surgery, a setting which is described in detail in Section 2, and woven throughout the work.
+In an effort to summarize temporal trends, estimates are projected onto candidate marginal structural models (MSM), each representing a different pattern in how effect estimates vary across treatment initiation time (e.g., constant, linear, etc.).
+An MSM selection procedure is then applied to choose which structure best describes temporal variation in effect estimates.
+Next, to disentangle changes in treatment effects from changes in the patient population receiving treatment, we define a causal estimand that corresponds to the treatment effect which would have been observed if the covariate distribution among eligible patients at a given time instead reflected that at some alternative time.
+Finally, we introduce a metric summarizing variability across these “cross-temporal” contrasts to quantify the role of covariate shift in explaining observed effect changes across calendar time.
+To our knowledge, this work is the first to apply tools from semiparametric theory (Tsiatis, 2006, Kennedy, 2023) to study calendar time-varying treatment effects in EHR-based studies.
+As mentioned, this work is directly motivated by the use of EHR data to study weight loss outcomes following bariatric surgery, particularly in comparison to similarly obese patients who do not undergo surgery, a setting which is described in detail in Section 2, and woven throughout the work.
 Section 3 formalizes notation and describes calendar time-specific causal estimands.
 Sections 4 and 5 introduce statistical tools, based on semiparametric efficiency theory, for determining how and why causal effects vary over calendar time, respectively, while Section 6 outlines how analysts can synthesize information from estimation and inferential procedures outlined in previous sections to determine what to report in their own study.
-Section 7 returns to the motivating study to apply the methods outlined in this work, and finally, Section 8 concludes with discussion.
+Section 7 presents results leveraging tools from this framework to detect and characterize calendar time-varying effects in our study of weight loss following bariatric surgery.
+Finally, Section 8 concludes with discussion.
 Detailed proofs and simulations which validate tools from this framework are provided in the Supplementary Material.
-2 An EHR-based Study of Bariatric Surgery
-Bariatric surgery is an intervention to mitigate obesity and related comorbidities, with typical candidates having a body mass index (BMI) exceeding 35 kg/m2 (kaiserpermanente2023).
-Due to the cost and complexity of conducting large-scale trials for bariatric surgery (Courcoulas2024), EHR-based studies are a critical tool to study long-term outcomes following bariatric surgery, particularly in relation to similarly obese patients not undergoing surgery (arterburn2020, obrien2018microvascular, fisher2018association, Harrington2024, coleman2022bariatric).
-Despite its success as a weight-loss intervention, only 1-2% of eligible candidates elect to undergo bariatric surgery annually (arterburn2014bariatric).
+2 An EHR-based Study of Weight Loss Following Bariatric Surgery
+2.1 Bariatric Surgery
+Bariatric surgery is an intervention to mitigate obesity and related comorbidities, with typical candidates having a body mass index (BMI) exceeding 35 kg/m2.
+Due to the cost and complexity of conducting large-scale trials for bariatric surgery (Courcoulas et al., 2024), EHR-based studies are a critical tool to study long-term outcomes following bariatric surgery, particularly in relation to similarly obese patients not undergoing surgery (Arterburn et al., 2020, O’Brien et al., 2018, Fisher et al., 2018, Harrington et al., 2024, Coleman et al., 2022).
+Despite its success as a weight-loss intervention, only 1-2% of eligible candidates elect to undergo bariatric surgery annually (Arterburn and Courcoulas, 2014).
 Therefore, EHR-based studies of bariatric surgery must consider a multi-year study entry period to obtain a meaningful number of patients undergoing surgery.
-The two most common bariatric surgical procedures are Roux-en-Y gastric bypass (RYGB) and sleeve gastrectomy (SG), with SG having surpassed RYGB in frequency in the past 15 years (arterburn2014bariatric, Reames2014).
+The two most common bariatric surgical procedures are Roux-en-Y gastric bypass (RYGB) and sleeve gastrectomy (SG), with SG having surpassed RYGB in frequency in the past 15 years (Arterburn and Courcoulas, 2014).
 SG is a less invasive and technically easier procedure than RYGB, and as such is offered by more surgeons.
-Relevant to this work is that the surgical technique for SG has historically been somewhat less standardized than for RYGB, particularly during the time-frame during which SG became the more commonly performed procedure (Guldogan2020Sleeve, Rosenthal2012, Gagner2013).
-Given the rapid shifts in the relative popularity of RYGB and SG in recent years, along with possibly evolving surgical techniques and standards of care for obesity management, understanding potential temporal changes in the effectiveness bariatric surgery is of particular interest.
+Relevant to this work is that the surgical technique for SG has historically been somewhat less standardized than for RYGB, particularly during the time-frame during which SG became the more commonly performed procedure (Guldogan et al., 2020, Rosenthal et al., 2012, Gagner et al., 2013).
+Given the rapid shifts in the relative popularity of RYGB and SG in recent years, along with possibly evolving surgical techniques and standards of care for obesity management, understanding potential temporal changes in the effectiveness of bariatric surgery is of particular interest.
+Not only does characterization of this temporal structure offer insight into these procedures during an important change in their utilization history, but it also includes assessments of treatment effects at the end of the study period that are more relevant to current clinical practice when the data are proximal to present day.
+2.2 The DURABLE EHR Database
 With this backdrop, we examined relative weight change at four pre-specified times post-surgery (6 months, 1 year, 2 years, 3 years), comparing 17,905 patients undergoing bariatric surgery between 2005-2011 at one of three Kaiser Permanente (KP) sites (KP Washington, KP Northern California, KP Southern California) with 933,044 non-surgical patients.
 In particular, we compared outcomes between surgery and no surgery, each of RYGB/SG versus no surgery separately, and RYGB versus SG.
-We leveraged EHR data from DURABLE, an NIH-funded study examining long-term outcomes of bariatric surgery across Kaiser Permanente sites (arterburn2020, coleman2016longterm, obrien2018microvascular, coleman2022bariatric, fisher2018association, Harrington2024, haneuse2025).
-Using this data source, arterburn2020 previously characterized longitudinal weight trajectories following RYGB and SG in comparison to matched non-surgical patients, though calendar time was considered only as a potential confounder.
-As an alternative to the matched cohort design employed by arterburn2020, we used a sequential target trial emulation approach (benz2024, madenci2024, haneuse2025, cain2010, danaei2013observational, caniglia2023).
+We leveraged EHR data from DURABLE, an NIH-funded study examining long-term outcomes of bariatric surgery across Kaiser Permanente sites (Arterburn et al., 2020, Coleman et al., 2016, O’Brien et al., 2018, Coleman et al., 2022, Fisher et al., 2018, Harrington et al., 2024, Haneuse et al., 2025).
+Using this data source, Arterburn et al.
+Arterburn et al.
+(2020) previously characterized longitudinal weight trajectories following RYGB and SG in comparison to matched non-surgical patients, though calendar time was considered only as a potential confounder.
+As an alternative to the matched cohort design employed by Arterburn et al.
+Arterburn et al.
+(2020), we used a sequential target trial emulation approach (Benz et al., 2025a, Madenci et al., 2024, Haneuse et al., 2025, Danaei et al., 2013, Caniglia et al., 2023).
 Specifically, we considered a sequence of 84 target trials (formalized in Section 3), initiating at monthly intervals between January 2005 and December 2011.
-Following arterburn2020, our study population included adults aged 19-79 years with BMI 35 kg/m2.
+Following Arterburn et al., our study population included adults aged 19-79 years with BMI exceeding 35 kg/m2.
 We excluded patients without a year of continuous enrollment in KP insurance and those pregnant in the year prior to the start of a particular target trial.
-17,596 surgical patients (98.2%) and 763,016 non-surgical patients (81.8%) met the study inclusion criteria at baseline of at least one target trial.
+In total, 17,596 surgical patients (98.2%) and 763,016 non-surgical patients (81.8%) met the study inclusion criteria at baseline of at least one target trial.
 Figure 1: Distribution of patient characteristics in DURABLE electronic health record database over calendar time for eligible patients undergoing bariatric surgery between 2005-2011.
@@
 $m=1$ corresponds to January 2005 and $m=84$ corresponds to December 2011.
-On the basis of these data, one might choose to estimate and report common a effect for bariatric surgery on 3 year relative weight change, in comparison to non-surgical patients.
-Though formal details are deferred until Section 7, such an approach produced estimates of -19.8% (-19.9%, -19.6%) for surgery overall, with corresponding estimates of -23.1% (-23.3%, -22.9%) for RYGB and -16.3% (-16.6%, -16.1%) and SG, in close agreement with estimates of -23.5% and -16.7% reported by arterburn2020, respectively.
-These summaries, however, do not acknowledge the potential for variation across calendar time, and thus do not offer any insight into possible temporal variation in the effectiveness of bariatric surgery.
-Figure 1 summarizes the distribution of several key covariates among eligible bariatric patients across the 84 target trials (i.e., across calendar time).
-Notably, the distribution of several effect modifiers highlighted by arterburn2020 (age, BMI, race, and diabetes status) varies meaningfully across time of surgery.
-For example, the mean BMI of patients undergoing bariatric surgery in a given calendar month declines from 47.5 kg/m2 in January 2005 to 43.8 kg/m2 in December 2011, with monthly averages ranging between 42.9 kg/m2 and 49.6 kg/m2.
+2.3 Calendar Time-Varying Effects in Bariatric Surgery
+Towards quantifying the relative benefits of surgery versus no surgery, the standard way forward would be to estimate and report a common effect across calendar time, for example, on 3-year relative weight change.
+Though formal details are deferred until Section 7, such an approach produces estimates of -19.8% (-19.9%, -19.6%) for surgery overall, with corresponding estimates of -23.1% (-23.3%, -22.9%) for RYGB and -16.3% (-16.6%, -16.1%) for SG.
+Note, these are in close agreement with estimates of -23.5% and -16.7% reported by Arterburn et al.
+Arterburn et al.
+(2020), respectively.
+While such summaries have appeal, particularly for communication to broad audiences, they may conceal meaningful changes in the effectiveness of bariatric surgery over calendar time, whether due to evolving surgical techniques, shifting patient populations, or perhaps both.
+Towards gaining preliminary insight, Figure 1 summarizes the distribution of several key covariates among eligible bariatric patients across the 84 target trials (i.e., across calendar time).
+Notably, the distribution of several effect modifiers highlighted by Arterburn et al.
+Arterburn et al.
+(2020) (age, BMI, race, and diabetes status) varies meaningfully across calendar time.
+For example, mean BMI of patients undergoing bariatric surgery in a given calendar month declines from 47.5 kg/m2 in January 2005 to 43.8 kg/m2 in December 2011, with monthly averages ranging between 42.9 kg/m2 and 49.6 kg/m2.
 Accordingly, any attempt to study how the effect of bariatric surgery may have changed over the course of the study period should acknowledge and account for this covariate shift.
@@
 3 Problem Setting
+In a hypothetical clinical trial, each patient would be randomized to undergo or not undergo surgery and be followed from the point of treatment assignment.
+By contrast, patients in our EHR-based study are not randomized to their respective treatment assignment.
+A key challenge in addressing the primary questions laid out in Section 2 is that, in contrast to patients who undergo surgery, there is no obvious way to assign a “time zero” to patients who are in the “no surgery” arm (Hernán et al., 2016).
+To resolve this dilemma, we partition the study entry period into a sequence of $M$ target trials (Benz et al., 2025a, Madenci et al., 2024, Danaei et al., 2013, Caniglia et al., 2023), indexed by $m\in\{1,\dots,M\}$.
+At baseline of each target trial, study eligibility is queried for all patients in the EHR, and patients are included in each trial for which they meet the eligibility criteria.
+In our study of bariatric surgery, $M=84$ and component target trials are conducted at monthly intervals between January 2005 ($m=1$) and December 2011 ($m=M=84$).
 3.1 Notation
-When a treatment strategy does not correspond to a concrete clinical event, time zero can be challenging to define (hernan2016immortal).
-In order to ensure that the start of follow-up is well-defined for all patients, particularly those not undergoing surgery, we partition the study entry period into a sequence of $M$ target trials (benz2024, madenci2024, cain2010, danaei2013observational, caniglia2023), indexed by $m\in\{1,\dots,M\}$.
-In the motivating study of bariatric surgery, $M=84$ and component target trials are conducted at monthly intervals between January 2005 ($m=1$) and December 2011 ($m=M=84$).
-To the best of our knowledge, the sequential target trial emulation design has been applied exclusively in applications examining time-to-event outcomes (scola2023implementation).
-Since relative weight change is not a time-to-event endpoint, we describe notation to formalize this design for continuous outcomes.
-Let $A_{m}\in\{0,1\}$ denote a binary treatment (e.g., surgery vs.
+To the best of our knowledge, the sequential target trial emulation design has been applied exclusively in applications examining time-to-event outcomes (Scola et al., 2023).
+Since relative weight change is not a time-to-event endpoint, we describe notation to formalize this design for univariate continuous outcomes.
+Let $A_{m}\in\{0,1\}$ denote a binary point exposure (e.g., surgery vs.
 no surgery) received at baseline of trial $m$, $Y_{m}\in\mathbb{R}$ denote an outcome of interest (e.g., relative weight change since baseline of trial $m$), and $\bm{L}_{m}\in\mathbb{R}^{p}$ denote a set of covariates at baseline of trial $m$ assumed sufficient for control of confounding, such as those in Figure 1.
-Let $E_{m}=g(\bm{L}_{m})$ be a binary eligibility indicator for inclusion in target trial $m$, where $g(\cdot)$ represents a fixed and known eligibility rule (benz2024, benz2025).
+Let $E_{m}=g(\bm{L}_{m})$ be a binary eligibility indicator for inclusion in target trial $m$, where $g(\cdot)$ represents a fixed and known eligibility rule (Benz et al., 2025a, b).
 Practically, because not all patients are in the EHR for the entire duration of the study entry period, we treat subjects as ineligible $(E_{m}=0)$ for periods before they appear and/or after they disappear from the EHR.
-Thus, at each of $M$ regularly occurring intervals, we ascertain $(\bm{L}_{m},A_{m},Y_{m})$ for eligible subject-trials; formally, we work with coarsened observed data units:
-$O=(E_{1},E_{1}\bm{L}_{1},E_{1}A_{1},E_{1}Y_{1},\dots,E_{m},E_{m}\bm{L}_{m},E_{m}A_{m},E_{m}Y_{m},\dots,E_{M},E_{M}\bm{L}_{M},E_{M}A_{M},E_{M}Y_{M}).$
+Thus, at each of $M$ regularly occurring intervals, we ascertain for each patient $i$ their eligibility $E_{m,i}$ for trial $m$, and, if eligible, $\bm{Z}_{m,i}\coloneqq(\bm{L}_{m,i},A_{m,i},Y_{m,i})$.
+That is, patients contribute $\bm{Z}_{m,i}$ for all target trials where they meet study eligibility.
+Formally, we work with coarsened observed data units
+$O_{i}=(E_{1,i},E_{1,i}\bm{Z}_{1,i},\dots E_{m,i},E_{m,i}\bm{Z}_{m,i},\dots E_{M,i},E_{M,i}\bm{Z}_{M,i})$
 We observe a sample $(O_{1},\ldots,O_{n})$ of $n$ independent and identically distributed observations from an underlying and unknown data distribution, $P$, i.e., the joint distribution of information across both patients and trials.
-Implicitly throughout, we assume covariates across trials share a common domain $\mathcal{L}$.
-As will become apparent in Sections 3.2 and 4, covariate, treatment and outcome values of ineligible patients are not required for the definition of our calendar time-specific causal estimands, nor for estimation.
+For succinctness, we often drop the patient index $i$ and work with a generic observation $O$.
+Implicitly throughout, we assume covariates across trials share a common support contained in $\mathcal{L}$.
+As will become clear in Sections 3.2 and 4, covariate, treatment and outcome values of ineligible patients are not required for the definition of our calendar time-specific causal estimands, nor for estimation.
 For any $P$-integrable function $f$ we define $Pf=\mathbb{E}_{P}[f]=\int f\,dP$ to denote the expectation of $f$ under $P$, and analogously define $\mathbb{P}_{n}f=\frac{1}{n}\sum_{i=1}^{n}f(O_{i})$ to express the sample mean of $f$ on data units $(O_{1},...,O_{n})$.
-Lastly, we let $Y_{m}(a_{m})$ denote counterfactual outcomes at time $m$, that is the outcome that would have been observed under treatment $A_{m}=a_{m}$ (hernan2024).
+Lastly, we let $Y_{m}(a_{m})$ denote counterfactual outcomes for the $m^{\text{th}}$ trial, that is the outcome that would have been observed under treatment $A_{m}=a_{m}$ (Hernán and Robins, 2024).
 3.2 Calendar Time-Specific Average Treatment Effects
@@
 Implicit in the definition of $\chi_{m}(P)$ is that comparisons are being made between the specific versions of treatment (i.e., surgical technique, standard of care) in place at time $m$.
-Note, $\chi_{m}(P)$ is an intention-to-treat (ITT) effect (hernan2016target) in the sense that it captures the effect of initiating treatment at time $m$ $(a_{m}=1)$ versus not initiating $(a_{m}=0)$ without regard to what happens at future time points.
-In particular, the definition of $\chi_{m}(P)$ is not concerned that treatment definitions may evolve over time (e.g., $a_{j}=0$ and $a_{m}=0$ may mean slightly different things for $j>m$), nor for the possibility of treatment switching.
+We are not interested in defining nor estimating effects of dynamic treatment regimes and note that consistent with the definition of $A_{m}$ as a point exposure, $\chi_{m}(P)$ captures the effect of initiating treatment at time $m$ $(a_{m}=1)$ versus not initiating $(a_{m}=0)$ without regard to what happens at future time points.
+In particular, the definition of $\chi_{m}(P)$ is not concerned that treatment definitions may evolve over time (e.g., $a_{j}=0$ and $a_{m}=0$ may mean slightly different things for $j>m$), nor for the possibility of treatment switching; see Section 8 for a discussion of how our framing relates to the consistency assumption and “multiple versions of treatment”.
+Due to the rarity of initiating surgery relative to standard of care and since bariatric surgery is effectively an irreversible procedure (at least in so far as it relates to a patient adhering to surgery or standard of care), non-adherence is extremely uncommon in our study.
+Just as a constant hazard ratio represents a weighted average of time-varying hazard ratios (across time since treatment initiation) (Hernán, 2010), a sequential target trial emulation estimating a constant ATE across calendar time implicitly targets some weighted average of trial-specific effects $\chi_{m}(P)$.
+However, the precise definition and causal interpretation of such contrast is rarely formalized or made transparent.
 4 Estimation of Calendar Time-Specific Average Treatment Effects
+In order to answer the clinical questions raised in Section 2, one might initially target the nonparametric curve composed of the collection $\bigr\{\chi_{m}(P)\}_{m=1}^{M}$.
+As we elaborate upon, however, a small number of surgical patients in any particular target trial presents a challenge towards efficient estimation of any individual $\chi_{m}(P)$.
+MSMs have proven useful for analysis of continuous exposures by parameterizing a dose-response curve to make causal inference tractable Robins et al.
+(2000), Naimi et al.
+(2014), Hernán and Robins (2024).
+In that same vein, pragmatic adoption of MSMs in the present context provides structure on top of $\bigr\{\chi_{m}(P)\}_{m=1}^{M}$, allowing information to be borrowed across target trials to describe trends over calendar time.
+4.1 Assumptions
 Given that treatment decisions in EHR-based settings are not randomized, we begin by articulating standard causal inference assumptions to address the potential for confounding.
@@
 $Y_{m}(a_{m})\perp\!\!\!\perp A_{m}\penalty 10000\ |\penalty 10000\ \bm{L}_{m},E_{m}=1$ for $a_{m}\in\{0,1\}$ and all $m$.
-In our data application, we take $\bm{L}_{m}$ to be the most recent values of covariates outlined in Figure 1, although we note that analysts can also include in this vector appropriate summaries of covariate and/or treatment history prior to baseline of trial $m$, if warranted in their application.
+Following closely to Arterburn et al.
+(Arterburn et al., 2020), we take $\bm{L}_{m}$ to be the most recent values of covariates outlined in Figure 1, although we note that analysts can also include in this vector appropriate summaries of covariate and/or treatment history prior to baseline of trial $m$, if warranted in their application.
+A careful discussion of Assumption 1, particularly in the context of multiple versions of treatment and changing treatment efficacy, is deferred until Section 8.
 4.2 Motivating A Projection Parameter Approach
-At the outset, estimation of each calendar time-specific ATE $\chi_{m}(P)$ may seem straightforward.
-Let $\mu_{m}(a_{m},\bm{L}_{m})=\mathbb{E}[Y_{m}\penalty 10000\ |\penalty 10000\ A_{m}=a_{m},\bm{L}_{m},E_{m}=1]$ and $\pi_{m}(L_{m})=P(A_{m}=1\penalty 10000\ |\penalty 10000\ \bm{L}_{m},E_{m}=1)$ denote the trial $m$ specific outcome regression and propensity score models, respectively.
-Under Assumptions 1-3, $\chi_{m}(P)$ is identified by the $g$-formula, $\mathbb{E}[\mu_{m}(1,\bm{L}_{m})-\mu_{m}(0,\bm{L}_{m})\penalty 10000\ |\penalty 10000\ E_{m}=1]$ or via inverse probability weighting, $\mathbb{E}\big[\big(\frac{A_{m}}{\pi_{m}(\bm{L}_{m})}-\frac{1-A_{m}}{1-\pi_{m}(\bm{L}_{m})}\big)Y_{m}\penalty 10000\ |\penalty 10000\ E_{m}=1\big]$ (hernan2024).
-One might then proceed to estimate $\chi_{m}(P)$ on the basis of its nonparametric influence function (tsiatis_semiparametric_2006, kennedy2023semiparametric),
+Let $\mu_{m}(a_{m},\bm{L}_{m})=\mathbb{E}[Y_{m}\penalty 10000\ |\penalty 10000\ A_{m}=a_{m},\bm{L}_{m},E_{m}=1]$ and $\pi_{m}(L_{m})=P(A_{m}=1\penalty 10000\ |\penalty 10000\ \bm{L}_{m},E_{m}=1)$ denote the trial $m$-specific outcome regression and propensity score models, respectively.
+Under Assumptions 1-3, $\chi_{m}(P)$ is identified by the $g$-formula, $\mathbb{E}[\mu_{m}(1,\bm{L}_{m})-\mu_{m}(0,\bm{L}_{m})\penalty 10000\ |\penalty 10000\ E_{m}=1]$ or via inverse probability weighting, $\mathbb{E}\big[\big(\frac{A_{m}}{\pi_{m}(\bm{L}_{m})}-\frac{1-A_{m}}{1-\pi_{m}(\bm{L}_{m})}\big)Y_{m}\penalty 10000\ |\penalty 10000\ E_{m}=1\big]$ (Hernán and Robins, 2024).
+One might then proceed to estimate $\chi_{m}(P)$ on the basis of its nonparametric influence function (Tsiatis, 2006, Kennedy, 2023),
 $\dot{\chi}^{*}_{m}(O;P)=\frac{\mathds{1}(E_{m}=1)}{P(E_{m}=1)}$ $\biggr\{\mu_{m}(1,\bm{L}_{m})-\mu_{m}(0,\bm{L}_{m})-\chi_{m}(P)$
@@
 Unfortunately, estimating $\mu_{m}$ or $\pi_{m}$ by fitting a separate model for each trial would be difficult in practice, especially in settings where a relatively small number of patients initiate a particular treatment during trial $m$.
-For example, 24 of 84 target trials in our study have fewer than 100 patients undergoing bariatric surgery at the corresponding baseline month, and 12 of 84 of have fewer than 25 surgical patients.
-This problem is further compounded for comparisons involving the individual surgical procedures, RYGB and SG (Figure S2).
-The predominant analytic strategy in sequential target trial emulations for time-to-event outcomes involves fitting a single parametric outcome and/or propensity score model pooling information across trials (scola2023implementation) (e.g., pooled logistic regression (dagostino1990)), perhaps with a spline term to flexibly adjust for calendar time as a confounder (dickerman2019, madenci2024).
+For example, 24 of 84 target trials in our study have fewer than 100 patients undergoing bariatric surgery at the corresponding baseline month, and 12 of 84 have fewer than 25 surgical patients.
+In fact, the lack of power from small sample sizes inherent to emulating a target trial from a single enrollment window was one of the primary motivations behind development of the sequential target trial design (Hernán et al., 2008).
+The predominant analytic strategy in sequential target trial emulations for time-to-event outcomes targets (discrete time) hazard ratios by fitting a single parametric outcome and/or propensity score model pooling information across trials (Scola et al., 2023) (e.g., pooled logistic regression (D’Agostino et al., 1990)), perhaps with spline terms to enable non-proportional hazards and flexibly adjust for calendar time as a confounder (Dickerman et al., 2019, Madenci et al., 2024).
 Given that trends in $\chi_{m}(P)$ over calendar time may reflect a complex interplay between changing treatment efficacy and shifts in effect modifier distributions, applying a pooled parametric model in this setting would require correctly specifying a substantial number of flexible interactions between treatment, confounders, and calendar time.
-Such a task may be unrealistic in many instances—arguably including our data application given the heterogeneity reported in Figure 1—and might instead motivate the use of pooled models based on flexible machine learning methods.
-Ultimately, while we are interested in trial-specific ATEs, we are most interested in the underlying trend in $\chi_{m}(P)$ across calendar time.
-To formalize estimation of this trend across calendar time, we follow the strategy of kennedy2019 and others (kennedy2023density, neugebauer2007nonparametric, rosenblum2010targeted), and formulate our estimand as a projection of all $M$ trial-specific effects onto a candidate parametric model.
-To operationalize this approach, we first define a MSM for mean trial-specific counterfactual outcomes.
+Such a task may be unrealistic in many instances—arguably including our study of bariatric surgery given the heterogeneity reported in Figure 1—and might instead motivate the use of pooled models based on flexible machine learning methods.
+Ultimately, while we are interested in trial-specific ATEs, we are most interested in characterization of the underlying trend in $\chi_{m}(P)$ across calendar time.
+To formalize estimation of this trend across calendar time, we follow the strategy of Kennedy et al.
+Kennedy et al.
+(2019) and others (Kennedy et al., 2023, Neugebauer and Van der Laan, 2007, Rosenblum and Van der Laan, 2010), and formulate our estimand as a projection of all $M$ trial-specific effects onto a candidate parametric marginal structural model.
+Specifically, following the pragmatic use of MSMs for analysis of continuous exposures (Kennedy et al., 2019, Naimi et al., 2014, Hernán and Robins, 2024, Robins et al., 2000), we define a MSM for mean trial-specific counterfactual outcomes.
 $\mathbb{E}[Y_{m}(a_{m})\penalty 10000\ |\penalty 10000\ E_{m}=1]=\psi_{0}(m)+\psi(m)a_{m}.$ (1)
-When the MSM in Equation (1) is saturated, $\psi(m)=\chi_{m}(P)$.
-Instead, we adopt a parametric model for $\psi(m)$, denoted $\psi(m;\bm{\beta})$, and take our estimand to be the projection of $\chi_{m}(P)$ onto $\psi(m;\bm{\beta})$.
+When the MSM in Equation (1) is saturated, $\psi(m)=\chi_{m}(P)$, that is, we are targeting the true nonparametric effect curve across calendar time.
+Because such contrast is not easily estimable due to challenges described above, we adopt a working parametric model for $\psi(m)$, denoted $\psi(m;\bm{\beta})$, and take our estimand to be the projection of $\chi_{m}(P)$ onto $\psi(m;\bm{\beta})$.
+Despite being a working model, $\psi(m;\bm{\beta})$ remains a useful summary of how treatment efficacy has evolved over calendar time.
+Moreover, when study data are proximal to the present, the more recent end of this curve, $\psi(M;\bm{\beta})$, offers direct insight into treatment efficacy in current clinical practice.
 Formally, $\bm{\beta}(P)$ is the projection parameter defined by
-$\bm{\beta}(P)=\mathop{\mathrm{argmin}}_{\bm{\beta}\in\mathbb{R}^{k}}\sum_{m=1}^{M}h(m)\Bigr(\chi_{m}(P)-\psi(m;\bm{\beta})\Bigr)^{2},$
-where $h(m)$ is an optional trial-specific weight.
-Henceforth, we proceed with weights of the form $h(m)=w(m)P(E_{m}=1)$ to weight trial contributions according to their effective sample size, where $w(m)$ allows for optional weighting by other additional factors.
-This approach offers several advantages, which we elaborate on in subsequent sections.
-As will become clear in Section 4.2, estimation of $\chi_{m}(P)$ is a necessary step towards estimating $\psi(m;\bm{\beta})$, a task for which we advocate the use of pooled nuisance models based on flexible machine learning methods.
-Nevertheless, our projection parameter approach allows us to identify meaningful underlying trends in $\chi_{m}(P)$ even if some trial-specific effects are estimated imprecisely (e.g., due to small sample sizes).
-Although pooled nonparametric models for component nuisance functions will be an essential step in estimating $\chi_{m}(P)$, they provide limited insight into how information is shared across calendar time.
-Introducing the additional projection step makes this information sharing explicit and clarifies how effect estimates from individual trials are combined to learn trends over calendar time, thereby improving transparency relative to relying on pooled nonparametric models alone.
-Finally, this approach naturally supports a model selection procedure (described in Section 4.4) for comparing candidate MSMs to identify the functional form that best describes how treatment effects vary over calendar time, including the possibility of a single, constant effect.
+$\bm{\beta}(P)=\mathop{\mathrm{argmin}}_{\bm{\beta}\in\mathbb{R}^{k}}\sum_{m=1}^{M}w(m)P(E_{m}=1)\Bigr(\chi_{m}(P)-\psi(m;\bm{\beta})\Bigr)^{2},$
+where $P(E_{m}=1)$ weights trial contributions according to their effective sample size and $w(m)$ allows for optional weighting by other additional factors.
+We emphasize that $\psi(m;\bm{\beta})$ is not believed to be the true effect curve but rather a tractable summary of trends across calendar time.
+Because the form of $\psi(m;\bm{\beta})$ that best approximates the nonparametric curve $\bigr\{\chi_{m}(P)\}_{m=1}^{M}$ is unknown, we outline a model selection procedure (described in Section 4.5) for comparing candidate MSMs to identify the functional form that best describes how treatment effects vary over calendar time, including the possibility of a single, constant effect.
 4.3 Efficient Estimation of Projection Parameter
 We now introduce preliminary efficiency theory for the projection parameters $\bm{\beta}(P)$, which will motivate our proposed estimator.
-Since influence functions are by definition mean zero, it is helpful to work with uncentered versions, $\dot{\chi}_{m}(O;P)\coloneqq\frac{\mathds{1}(E_{m}=1)}{P(E_{m}=1)}\dot{\chi}^{\dagger}_{m}(O;P)$, which satisfy $\mathbb{E}_{P}[\dot{\chi}_{m}(O;P)]=\chi_{m}(P)$, where
+Since influence functions are by definition mean zero, it is helpful when constructing estimators to work with uncentered versions, $\dot{\chi}_{m}(O;P)\coloneqq\frac{\mathds{1}(E_{m}=1)}{P(E_{m}=1)}\dot{\chi}^{\dagger}_{m}(O;P)$, which satisfy $\mathbb{E}_{P}[\dot{\chi}_{m}(O;P)]=\chi_{m}(P)$, where
 $\dot{\chi}^{\dagger}_{m}(O;P)\coloneqq\mu_{m}(1,\bm{L}_{m})-\mu_{m}(0,\bm{L}_{m})+\biggr(\frac{A_{m}}{\pi_{m}(\bm{L}_{m})}-\frac{1-A_{m}}{1-\pi_{m}(\bm{L}_{m})}\biggr)\Bigr(Y_{m}-\mu_{m}(A_{m},\bm{L}_{m})\Bigr).$
-These quantities play an important role in the influence function of $\bm{\beta}(P)$, as summarized in the following theorem.
 Theorem 1
-Suppose that the candidate MSM, $\psi(m;\bm{\beta})$ is twice differentiable in $\bm{\beta}$. Then, under Assumptions 1-3, $\bm{\beta}(P)$ is identified as the solution to the estimating equation,
+Suppose that the candidate MSM, $\psi(m;\bm{\beta})$ is twice differentiable in $\bm{\beta}$. Under Assumptions 1-3, $\bm{\beta}(P)$ is identified as the solution to the estimating equation,
 $0=\sum_{m=1}^{M}w(m)P(E_{m}=1)\nabla_{\bm{\beta}}\psi(m;\bm{\beta})[\chi_{m}(P)-\psi(m;\bm{\beta})].$
@@
 Equations (2) and (3) make clear, however, that to even estimate $\bm{\beta}(P)$, estimates of $\chi_{m}(P)$, and thus estimates of nuisance functions $\mu_{m}$ and $\pi_{m}$, are required.
-Also of note in Equation (3), trials are weighted proportionally to the sample size of their eligible populations, potentially mitigating the influence of noisier estimates of $\chi_{m}(P)$ in component trials with smaller sample sizes.
 […]
-In order to estimate $\chi_{m}(P)$, we first create a pooled dataset $\mathcal{D}$ of patient information across all trials in which they are eligible, that is,
+To estimate $\chi_{m}(P)$, we first create a pooled dataset $\mathcal{D}$ of patient information across all trials in which they are eligible, that is,
 $\mathcal{D}=\bigcup_{m=1}^{M}\Bigr\{(\bm{L}_{m,i},A_{m,i},Y_{m,i},m)\Bigr\}_{i:E_{m,i}=1}.$ (4)
 […]
-Routines for estimation of influence function-based estimators like $\widehat{\chi}_{m}$ or $\widehat{\bm{\beta}}$ and typically involve sample splitting or cross-fitting, particularly when using nonparametric and/or machine learning techniques to estimate component nuisance functions (chernozhukov2018).
-We describe the complete procedure for estimating $\widehat{\chi}_{m}$ and $\widehat{\bm{\beta}}$ below.
+Routines for estimation of influence function-based estimators like $\widehat{\chi}_{m}$ or $\widehat{\bm{\beta}}$ typically involve sample splitting or cross-fitting, particularly when using nonparametric and/or machine learning techniques to estimate component nuisance functions (Chernozhukov et al., 2018).
+We describe the complete procedure for estimating $\widehat{\chi}_{m}$ and $\widehat{\bm{\beta}}$ in Algorithm 1 below.
 Algorithm 1 (Computation of $\widehat{\chi}_{m}$ and $\widehat{\bm{\beta}}$)
@@
   6. 6.
-    For a given candidate marginal structural model, $\psi(m;\bm{\beta})$, plug estimates $\{\widehat{\chi}_{m}\}_{m=1}^{M}$ into Equation (3) and solve for $\widehat{\bm{\beta}}$ (e.g., via nleqslv (R-nleqslv)).
+    For a given candidate marginal structural model, $\psi(m;\bm{\beta})$, plug estimates $\{\widehat{\chi}_{m}\}_{m=1}^{M}$ into Equation (3) and solve for $\widehat{\bm{\beta}}$ (e.g., via nleqslv (Hasselman, 2023)).
 4.4 Theoretical Properties of $\widehat{\bm{\beta}}$ and $\widehat{\chi}_{m}$
@@
 Additional commentary on Assumptions S1-S3 is provided in the Supplementary Materials.
-Pooled models can help ensure that the required rate conditions are met in component trials with smaller effective sample sizes.
+Practically, pooled models can help ensure that the required rate conditions are met in component trials with smaller effective sample sizes.
 As discussed in Section 4.3, however, pooled parametric models for $\pi_{m}$ and $\mu_{m}$ are likely to be inconsistent in the presence of significant covariate shift, motivating the use of more flexible pooled models.
@@
 […]
-4.4 Model Selection Among Candidate Marginal Structural Models
-Suppose one considers a range of candidate MSMs including a constant model (i.e., $\psi(m;\beta)=\beta$), a linear model (i.e., $\psi(m;\bm{\beta})=\beta_{0}+\beta_{1}m$), and more flexible MSMs including higher-order polynomials (in $m$) and splines.
-A model selection procedure to select the MSM which best describes the underlying trend $\chi_{m}(P)$ is important to best characterize how treatment effects—and thus possibly treatment efficacy—vary over calendar time.
-In particular, if a constant model were selected, it would suggest that any underlying variation in treatment effects is not detectable given the study sample size, and thus may not be clinically meaningful.
-On the other hand, if a flexible MSM were selected, it would provide evidence of non-constant trends in treatment effects over calendar time.
-To implement such a model selection procedure, we adapt the ideas of kennedy2019 and vanderLaanDudoit2003cv.
+4.5 Selection Among Candidate Marginal Structural Models
+Unless an analyst a priori desires to summarize treatment efficacy through a single constant effect across calendar time, it is often unclear how to specify the functional form for $\psi(m;\bm{\beta})$ which appropriately summarizes the nonparametric curve $\bigr\{\chi_{m}(P)\}_{m=1}^{M}$, and in practice several candidate models may be reasonable.
+Forcing a single specification risks either under- or over-smoothing the underlying trend in $\chi_{m}(P)$, obscuring clinically meaningful variation or introducing spurious structure.
+It is therefore preferable to consider a range of candidate MSMs, including a constant model (i.e., $\psi(m;\beta)=\beta$), a linear model (i.e., $\psi(m;\bm{\beta})=\beta_{0}+\beta_{1}m$), and more flexible MSMs including higher-order polynomials (in $m$) and splines, and select among them using a data-adaptive procedure that lets the data inform the appropriate level of complexity.
+To implement such an MSM selection procedure, we adapt the ideas of Kennedy et al.Kennedy et al.
+(2019) and van der Laan and Dudoit van der Laan and Dudoit (2003).
 Let $\{\widehat{\psi}_{1},...,\widehat{\psi}_{K}\}$ denote a set of $K$ candidate MSM projection models, with $\widehat{\psi}_{k}\equiv\psi_{k}(m;\widehat{\bm{\beta}}_{k})$.
 A natural quantity by which to compare candidate MSMs is mean-squared error (relative to the true $\chi_{m}(P)$).
-To simplify comparisons, we drop terms that don’t depend on the candidate MSM and proceed with the pseudorisk,
+To simplify comparisons, we drop terms that do not depend on the candidate MSM and proceed with the pseudorisk,
 $L(\widehat{\psi}_{k})=\sum_{m=1}^{M}w(m)\bigr\{\psi_{k}(m;\widehat{\bm{\beta}}_{k})^{2}-2\psi_{k}(m;\widehat{\bm{\beta}}_{k})\chi_{m}(P)\bigr\}.$
@@
 […]
-This is an intuitive loss function to work with because it reflects an estimating equation-based estimator of the pseudorisk, based on its influence function, treating $\widehat{\psi}_{k}$ as fixed (kennedy2019).
+This is an intuitive loss function to work with because it reflects an estimating equation-based estimator of the pseudorisk, based on its influence function, treating $\widehat{\psi}_{k}$ as fixed (Kennedy et al., 2019).
 […]
 Figure 2: Overview of cross-fitting procedure for evaluating $\widehat{L}(\widehat{\psi}_{k})$.
-In practice, differences between candidate projection models can be small (kennedy2023semiparametric), particularly if models classes are nested within one another.
-In such cases, we might have a preference for simpler models, such that more flexible MSMs are selected only if there are clinically meaningful trends in $\chi_{m}(P)$ across calendar time.
+In practice, differences between candidate projection models can be small (Kennedy, 2023), particularly if model classes are nested within one another.
+In such cases, analysts might have a preference for simpler models, such that more flexible MSMs are selected only if there are clinically meaningful trends in $\chi_{m}(P)$ across calendar time.
 To formalize this, let $\widehat{\psi}^{*}=\mathop{\mathrm{argmin}}_{\widehat{\psi}_{k}\in\{\widehat{\psi}_{1},\dots,\widehat{\psi}_{K}\}}\widehat{L}(\widehat{\psi}_{k})$ denote the MSM among our set of candidates which minimizes the loss function $\widehat{L}(\widehat{\psi}_{k})$.
@@
 Finally, one can choose the simplest model such that $|\widehat{L}(\widehat{\psi}_{k})-\widehat{L}(\widehat{\psi}^{*})|\leq c\epsilon$, for some user-selected constant $c$.
-There is considerable debate in the literature about how large an effect size needs to be on a standard deviation scale to be clinically meaningful (tsujimoto2022mic), though the notion of clinically meaningful is very application dependent.
-In simulations, we find that $c=1/4$ works well in our setup, but present result for a range of other values.
-Finally, we note that notion of “simpler model” can be operationalized through any model complexity criterion so long as models can be ordered by this criterion.
+There is considerable debate in the literature about how large an effect size needs to be on a standard deviation scale to be clinically meaningful (Tsujimoto et al., 2022), though the notion of clinically meaningful is very application dependent.
+In simulations, we find that $c=1/4$ works well in our setup, but we present results for a range of other values.
+Finally, we note that the notion of “simpler model” can be operationalized through any model complexity criterion so long as models can be ordered by this criterion.
 One reasonable such choice—which we adopt in our application—is ordering candidate MSMs by the dimension of their indexing projection parameter (i.e., the dimension of $\bm{\beta}$).
 5 Decomposing Sources of Variation
-While the projection parameter framework and corresponding model selection procedure outlined in Section 4 provide insight into how treatment effects vary across calendar time, they fail to explain why such changes (if any) are observed.
+While the projection parameter framework and corresponding MSM selection procedure outlined in Section 4 provide insight into how treatment effects vary across calendar time, they fail to explain why such changes (if any) are observed.
 In particular, if the candidate MSM selected is non-constant, suggesting that $\chi_{m}(P)$ varies over time, we would like to ascertain the degree to which underlying changes in treatment efficacy—as opposed to distribution shift in effect modifiers—are responsible.
 5.1 A Cross-Trial Contrast to Remove Distribution Shift
-To isolate why treatment effects might vary over calendar time, we begin by defining new contrasts by reweighting mean counterfactual outcomes across eligible populations at different points in time:
+To isolate why treatment effects might vary over calendar time, we begin by defining new contrasts by averaging reweighted conditional average treatment effects (CATE) Dahabreh et al.
+(2019), Cole and Stuart (2010) over covariate distributions reflecting trial eligible populations at different points in time:
 $\chi_{j,m}(P)\coloneqq\int_{\mathcal{L}}\mathbb{E}\bigr[Y_{m}(a_{m}=1)-Y_{m}(a_{m}=0)\penalty 10000\ |\penalty 10000\ \bm{L}_{m}=\bm{\ell},g(\bm{L}_{m})=1\bigr]dP_{\bm{L}_{j}|E_{j}=1}(\bm{\ell}\penalty 10000\ |\penalty 10000\ E_{j}=1).$
-The estimand $\chi_{j,m}(P)$ represents the difference in mean counterfactual outcomes at time $m$ in a population where trial baseline covariates are standardized to the covariate distribution of the eligible population at time $j$.
+The estimand $\chi_{j,m}(P)$ represents the difference in mean counterfactual outcomes among patients eligible for the $m^{\text{th}}$ trial in a population where trial baseline covariates are standardized to the covariate distribution of the eligible population at time $j$.
 First, we note that when $j=m$, this contrast reduces to trial-specific ATE defined in Section 3.2, that is, $\chi_{m,m}(P)=\chi_{m}(P)$.
@@
 We exploit this fact in Section 5.3 to introduce a metric for quantifying effect modification.
-In order to identify each $\chi_{j,m}(P)$, we introduce an additional assumption similar to those commonly used in transportability analyses (ung2025, dahabreh2019generalizing, dahabreh2020extending).
+In order to identify each $\chi_{j,m}(P)$, we introduce an additional assumption similar to those commonly used in transportability analyses (Dahabreh et al., 2019, Cole and Stuart, 2010).
 A 4 (Cross-Trial Overlap)
@@
 As such, fitting a separate model for each $\xi_{j,m}$ is computationally infeasible.
-Moreover, conditional density estimation is a challenging statistical problem (wasserman2006all).
-Because $\xi_{j,m}$ is a ratio of conditional densities, however, we can directly apply Bayes’ rule (sugiyama2012density, benz2025, diaz2023) and leverage pooled models, as with $\widetilde{\mu},\widetilde{\pi}$:
+Moreover, conditional density estimation is a challenging statistical problem (Wasserman, 2006).
+Because $\xi_{j,m}$ is a ratio of conditional densities, however, we can directly apply Bayes’ rule (Sugiyama et al., 2012, Benz et al., 2025b, Díaz et al., 2023) and leverage pooled models, as with $\widetilde{\mu},\widetilde{\pi}$:
 $\xi_{j,m}(\bm{\ell})=\frac{p(\bm{L}_{j}=\bm{\ell}\penalty 10000\ |\penalty 10000\ E_{j}=1)}{p(\bm{L}_{m}=\bm{\ell}\penalty 10000\ |\penalty 10000\ E_{m}=1)}=\frac{P(E_{m}=1)P(T=j\penalty 10000\ |\penalty 10000\ \bm{L}=\bm{\ell},E=1)}{P(E_{j}=1)P(T=m\penalty 10000\ |\penalty 10000\ \bm{L}=\bm{\ell},E=1)},$ (6)
-where $T$ denotes the trial membership indicator in full the pooled data set, $\mathcal{D}$.
+where $T$ denotes the trial membership indicator in the full pooled data set, $\mathcal{D}$.
 Thus, we have turned the task of conditional density estimation into modeling the trial membership indicator conditional on $\bm{L}$ across all eligible subject-trials.
@@
 5.3 Summarizing Changes in Effect Estimates
-Ultimately, our interest in $\chi_{j,m}(P)$ is as a stepping stone towards determining the degree to which changes in calendar time-specific treatment effects are explained by changes in effects modifiers compared to something more fundamental in terms of changes in underlying treatment efficacy.
+Ultimately, our interest in $\chi_{j,m}(P)$ is as a stepping stone towards determining the degree to which changes in calendar time-specific treatment effects are explained by changes in effect modifiers compared to something more fundamental in terms of changes in underlying treatment efficacy.
 We let $\bm{S}\in\mathbb{R}^{M\times M}$ denote the standardization matrix of all $M^{2}$ cross-trial effects, with $S_{jm}=\chi_{j,m}(P)$, and let $\widehat{\bm{S}}$ denote the estimated standardization matrix with entries $\widehat{\chi}_{j,m}$.
-To quantify sources of variation in cross-trials effects, we define the following variance-like quantities:
+To quantify sources of variation in cross-trial effects, we define the following variance-like quantities:
 $\sigma_{j}^{2}$ $\coloneqq\frac{1}{M-1}\sum_{m=1}^{M}\Bigr(\chi_{j}(P)-\chi_{j,m}(P)\Bigr)^{2},$
@@
 Moreover, as discussed in Section 6, rejection of $H_{0}^{(1)}$ in favor of $\theta<1$ may motivate different standardization strategies when the primary goal is to understand changes in treatment efficacy.
-In practice, $H_{0}:\theta\in\{0,1\}$ considers testing at both boundary points for $\theta$, somewhat similar to testing variance components in mixed-effects models (stram1994variance) and as such characterizing the asymptotic distribution of $\widehat{\theta}$ under the null would likely yield a challenging ratio involving mixtures of chi-squared distributions.
+In practice, $H_{0}:\theta\in\{0,1\}$ considers testing at both boundary points for $\theta$, somewhat similar to testing variance components in mixed-effects models (Stram and Lee, 1994) and as such characterizing the asymptotic distribution of $\widehat{\theta}$ under the null would likely yield a challenging ratio involving mixtures of chi-squared distributions.
 On the other hand, a full nonparametric bootstrap is computationally infeasible in our setting, given that the size of the pooled dataset $\mathcal{D}$ exceeds several million eligible subject-trials for comparisons involving non-surgical patients (Figure S2).
 Not only does re-estimation of nuisance functions across bootstrapped datasets pose a large computational burden, but so too does prediction of $\widehat{\mu}_{m}(a_{m},\bm{L}_{j})$ and $\widehat{\xi}_{j,m}(\bm{L}_{m})$ given that the number of predictions required is a $M$-fold larger than the size of $\mathcal{D}$.
-Instead, we use an asymptotic version of the parametric bootstrap (van2000asymptotic, EfronTibshirani1994) for $\bm{S}$.
+Instead, we use an asymptotic version of the parametric bootstrap (Van der Vaart, 2000) for $\bm{S}$.
 In particular, under the conditions of Theorem 4, $\sqrt{n}\bigr(\widehat{\bm{S}}-\bm{S})\overset{d}{\to}\mathcal{N}(0,\bm{\Sigma})$ where $\bm{\Sigma}\in\mathbb{R}^{M^{2}\times M^{2}}$ is the covariance matrix of $\bm{S}$, with entries $\bm{\Sigma}(s_{m_{1}j_{1}},s_{m_{2}j_{2}})=\text{Cov}[\dot{\chi}^{*}_{j_{1},m_{1}}(O;P),\dot{\chi}^{*}_{j_{2},m_{2}}(O;P)]$.
@@
 Unfortunately, due to the boundary null, bootstrapped replicates $\widehat{\theta}^{(1)},\dots,\widehat{\theta}^{(B)}$ will never be exactly 0 or 1 and thus 95% percentile-based confidence intervals would never reject $H_{0}$.
-We follow a similar strategy to that of CavaliereNielsenPedersenRahbek2022, and modify our test to allow for a small margin away from the boundary.
+We follow a similar strategy to that of Cavliere et al.
+Cavaliere et al.
+(2022), and modify our test to allow for a small margin away from the boundary.
 For some small $\delta>0$, we adopt the test $H_{0}^{(\delta)}:\theta\in[0,\delta]\cup[1-\delta,1]$ versus $H_{1}^{(\delta)}:\theta\in(\delta,1-\delta)$, and reject $H_{0}^{(\delta)}$ if a 95% percentile-based confidence interval is fully within $(\delta,1-\delta)$.
-One can interpret $\delta$ as the fraction of variability in entries in $\bm{S}$ for which relative changes in causal effects is deemed negligible.
+One can interpret $\delta$ as the fraction of variability in entries in $\bm{S}$ for which relative changes in causal effects are deemed negligible.
 The choice of $\delta$ is highly context dependent, as what constitutes clinically meaningful changes is application specific; in the present context of bariatric surgery, we adopt $\delta=0.05$.
@@
 6.1 Guidance Under No Violations of Assumptions 1-4
-To illustrate how model selection and hypothesis testing for $\theta$ can be used together to understand reporting of calendar time-varying effects, we consider two scenarios, one where a constant MSM is selected and a second where a non-constant MSM is selected.
+To illustrate how MSM selection and hypothesis testing for $\theta$ can be used together to understand reporting of calendar time-varying effects, we consider two scenarios, one where a constant MSM is selected and a second where a non-constant MSM is selected.
 First, when the selected MSM is a constant, it is reasonable to report a common effect for the study population, $\psi(m;\widehat{\bm{\beta}})=\widehat{\beta}$, regardless of the value of $\theta$.
@@
 Extensive simulations, tied closely to the motivating study of bariatric surgery, which validate the utility of the tools outlined in Sections 4 and 5, are provided in Appendix S3.
-Simulations highlight how MSM model selection varies by sample size and choice of shrinkage parameter $c$, along with the ability of the inferential procedure described in 5.5 to distinguish between effect modification from underlying variation in treatment efficacy.
-7 Application of Framework to Bariatric Surgery
-Finally, we return to our motivating study of bariatric surgery and apply the proposed methods to examine trends over calendar time from 2005 to 2011.
+Simulations highlight how MSM selection varies by sample size and choice of shrinkage parameter $c$, along with the ability of the inferential procedure described in Section 5.5 to distinguish between effect modification from underlying variation in treatment efficacy.
+7 Evaluating Temporal Trends in Causal Effects of Bariatric Surgery
+Finally, we return to our motivating study of bariatric surgery and leverage the proposed methods to examine trends over calendar time from 2005 to 2011.
 Analyses were conducted across 16 settings, defined by four post-surgical follow-up times (6 months, 1 year, 2 years, and 3 years) and four treatment comparisons (no surgery versus surgery, no surgery versus RYGB or SG separately, and RYGB versus SG).
-No SG patients met study eligibility criteria until month $m=30$, so analysis for comparisons involving SG was restricted to trials $m\in\{30,\dots 84\}$.
-Random forests were used to estimate all pooled nuisance models, separately for each setting, adjusting for age, BMI, sex, race, KP study site, self-reported smoking status, hypertension, dyslipdeimia, type II diabetes, and prescriptions for antilipemic or antihypertensive medications, or insulin.
-This set of potential confounders directly followed from those used by arterburn2020.
+No SG patients met study eligibility criteria until month $m=30$, so analysis for comparisons involving SG was restricted to trials $m\in\{30,\dots,84\}$.
+Random forests were used to estimate all pooled nuisance models, separately for each setting, adjusting for age, BMI, sex, race, KP study site, self-reported smoking status, hypertension, dyslipidemia, type II diabetes, and prescriptions for antilipemic or antihypertensive medications, or insulin.
+This set of potential confounders directly followed from those used by Arterburn et al.
+Arterburn et al.
+(2020).
 Candidate MSMs for projection included constant, linear, and cubic models, and splines with 2 and 3 internal knots.
@@
 […]
-Based on selected MSMs, surgical patients experienced excess 3-year weight change (relative to non-surgical patients) of -27.4% (-30.1%, -24.8%) at $m=1$ and -18.3% (-18.8%, -17.8%) at $m=84$, reflecting a change of 9.1% estimated treatment effect between 2005 and 2011.
-This temporal shift indicates that the common effect of –19.6% reported in Section 2, obtained by projection onto a constant MSM, obscures an important aspect understanding the efficacy of bariatric surgery.
+Based on selected MSMs, surgical patients experienced excess 3-year weight change (relative to non-surgical patients) of -27.4% (-30.1%, -24.8%) at $m=1$ and -18.3% (-18.8%, -17.8%) at $m=84$, reflecting a change of 9.1% in the estimated treatment effect between 2005 and 2011.
+This temporal shift indicates that the common effect of –19.6% reported in Section 2, obtained by projection onto a constant MSM, obscures an important aspect of understanding the efficacy of bariatric surgery.
 Similar changes were observed for RYGB patients, with estimated 3-year weight change differences ranging from -28.3% (-31.3%, -25.3%) at $m=1$ to -23.0% (-23.5%, -22.5%) at $m=84$.
 As reflected in Figure 3, there is little change in the 3-year RYGB effect estimate after $m=20$.
-Unsurprisingly then, the common effect estimate of -23.1% reported in Section 2 is comprises a reasonable summary of the treatment effect for a much greater portion of the study period.
+Unsurprisingly then, the common effect estimate of -23.1% reported in Section 2 comprises a reasonable summary of the treatment effect for a much greater portion of the study period.
 A constant MSM was selected for the 3-year weight change outcome comparing SG patients relative to non-surgical patients.
@@
 RYGB vs. SG 3 Years Constant Constant (0.688, 0.793)
-Table 2: Summary of model selection results and 95% bootstrapped intervals for $\theta$.
+Table 2: Summary of MSM selection results and 95% bootstrapped intervals for $\theta$.
 The minimizing MSM denotes $\widehat{\psi}^{*}$ minimizing $\widehat{L}(\widehat{\psi}$), while the selected MSM denotes the simplest model within $c=0.25$ weighted standard deviations of the minimizer.
@@
 Nevertheless, that our method is able to discover this well-explained trend offers evidence of its utility.
-When making comparisons to non-surgical patients, much of the clinical literature aggregates all surgical procedures under a single label “bariatric surgery” (obrien2018microvascular, Harrington2024, benz2024, madenci2024, haneuse2025, fisher2018association).
+When making comparisons to non-surgical patients, much of the clinical literature aggregates all surgical procedures under a single label “bariatric surgery” (O’Brien et al., 2018, Harrington et al., 2024, Benz et al., 2025a, Madenci et al., 2024, Haneuse et al., 2025, Fisher et al., 2018).
 As such, recognizing that both the meaning and the underlying effectiveness of this aggregated treatment label have changed is particularly relevant.
-Our results also hint at possible improvements in SG technique and improved weight loss outcomes 6-months post surgery, though such improvements are not observed for longer-term outcomes.
-One possible explanation is that any underlying changes in technique have helped patients reach their weight-loss nadir more quickly, but have not influenced peak weight-loss or rebound effects (courcoulas2013weight, arterburn2020).
+Importantly, while the $m=84$ estimate of -18.3% is more reflective of current real-world efficacy of bariatric surgery than the common effect of -19.6%, differences are attributable to both changing distribution of procedures and patient populations being treated.
+As such additional standardization efforts may be required to get a truly contemporary estimate, as the DURABLE database only follows patients through 2015.
+Finally, results also hint at possible improvements in SG technique and improved weight loss outcomes 6 months post-surgery, though such improvements are not observed for longer-term outcomes.
+One possible explanation is that any underlying changes in technique have helped patients reach their weight-loss nadir more quickly, but have not influenced peak weight-loss or rebound effects.
 8 Discussion
@@
 The proposed methods generalize to other study designs, however, provided that the study entry period can be discretized into sufficiently narrow windows through which one can define calendar time-specific effects.
-For example, head-to-head comparisons of RYGB and SG do not strictly require a sequence of target trials because the start of follow-up is defined by the date of surgery, and a (matched) cohort design may be a reasonable alternative (McTigue2020, benz2025).
+For example, head-to-head comparisons of RYGB and SG do not strictly require a sequence of target trials because the start of follow-up is defined by the date of surgery, and a (matched) cohort design may be a reasonable alternative (McTigue et al., 2020, Benz et al., 2025b).
 Under such designs, one might partition the study entry period into small intervals and define contrasts analogous to $\chi_{m}(P)$ and $\chi_{j,m}(P)$, before applying the ideas introduced in this work.
 The choice of interval length will be application specific, and depend on both how quickly clinical practice can change as well as the frequency of treatment.
+The decomposition framework introduced in Section 5 shares conceptual motivation with Gilbert et al.
+(Gilbert et al., 2025), who propose a related decomposition to explain heterogeneity in treatment effects across distinct studies or sites.
+However, the two frameworks differ in important ways.
+Gilbert et al.
+decompose between-study heterogeneity into contributions from case-mix differences, mediating variables, and pure effect modification across externally defined, unordered study indices.
+By contrast our framework is specifically designed for the temporally ordered, calendar time setting, where the goal is to disentangle changes in treatment efficacy from covariate shift occurring within a single EHR-based study.
+Additionally, rather than estimating decomposition components separately, we introduce a scalar summary metric $\theta$ to quantify the aggregate role of covariate shift in explaining observed temporal variation in treatment effects.
 In studies where there is no clinical rationale for changes in underlying treatment efficacy (e.g., when treatments are fixed-dose pharmaceutical medications), summarizing variation in cross-trial effects can serve as a useful diagnostic for detecting effect modification over time.
 For instance, $\theta$ may be particularly informative in pharmacoepidemiologic studies, helping to quantify how treatment effects vary as the characteristics of medication users evolve with broader adoption.
-We note that there are natural connections between our work and prior work on violations of the consistency assumption in causal inference (cole2009consistency, vanderweele2009consistency, VanderWeele2013).
+We note that there are natural connections between our work and prior work on violations of the consistency assumption in causal inference (Cole and Frangakis, 2009, VanderWeele, 2009, VanderWeele and Hernán, 2013).
 While RYGB and SG can be framed as multiple versions of bariatric surgery, it is perhaps less natural to view each procedure over time as multiple versions of treatment in the usual sense, in part because it is rarely known when treatments differ, and thus which “version” of treatment was received.
 In a sense, our setup implicitly defines trial-specific versions of treatment, i.e., through the symbols $A_{m}$ and $a_{m}$.
-However, we choose to view counterfactual outcomes $Y_{m}(a_{m})$ similarly to those defined and framed in cole2009consistency, $Y(a,k)$, where the additional treatment details $(k)$ required to maintain consistency are entirely captured by trial index ($m$) in our notation.
-In particular, we adopt the viewpoint that if $a_{m}=a_{j}$ but $Y_{m}(a_{m})\neq Y_{j}(a_{j})$, then this reflects changes in the efficacy of a particularly treatment $a$ between two time points $m,j$, rather than always calling this phenomenon multiple versions of treatment.
-This choice is in line with Cole and Frangakis, who note that taking versions of treatment to the extreme precludes any sort of generalizability and hampers the utility of a study.
+However, we choose to view counterfactual outcomes $Y_{m}(a_{m})$ similarly to those defined and framed in Cole and Frangakis Cole and Frangakis (2009), $Y(a,k)$, where the additional treatment details $(k)$ required to maintain consistency are entirely captured by trial index ($m$) in our notation.
+In particular, we adopt the viewpoint that if $a_{m}=a_{j}$ but $Y_{m}(a_{m})\neq Y_{j}(a_{j})$, then this reflects changes in the efficacy of a particular treatment $a$ between two time points $m,j$, rather than always calling this phenomenon multiple versions of treatment.
+This choice is in line with Cole and Frangakis, who note that taking versions of treatment to the extreme precludes generalizability and hampers the utility of a study.
+For instance, we would find it counterproductive to define separate versions of bariatric surgery for each operating surgeon, or of non-surgical weight management based on prevailing pharmacological or diet and exercise standards.
 Finally, we have assumed throughout that during the time period where a patient is enrolled in Kaiser Permanente insurance, and thus is accessible in the EHR, there is no missing data in components of $O$.
@@
 For the purposes of this work, we treat patients with incomplete information at a particular trial as ineligible $(E_{m}=0)$.
-However, we see combining ideas from our framework with recent work on causal inference under missing eligibility (benz2024, benz2025) as key area of future research.
+However, we see combining ideas from our framework with recent work on causal inference under missing eligibility (Benz et al., 2025a, b) as a key area of future research.
 Simultaneously handling possibly missing treatment status and/or outcomes across trials represents another important line of further methodological development.